### v1

# SplineSketch: Even More Accurate Quantiles with Error Guarantees

Abstract.
Space-efficient estimation of quantiles in massive datasets is a fundamental problem with numerous applications in data monitoring and analysis. While theoretical research led to optimal algorithms, such as the Greenwald-Khanna algorithm or the KLL sketch, practitioners often use other sketches that perform significantly better in practice but lack theoretical guarantees. Most notably, the widely used $t$-digest has unbounded worst-case error.
In this paper, we seek to get the best of both worlds. We present a new quantile summary, SplineSketch, for numeric data, offering near-optimal theoretical guarantees and outperforming $t$-digest by a factor of 2-20 on a range of synthetic and real-world datasets with non-skewed frequency distributions. To achieve such performance, we develop a novel approach that maintains a dynamic subdivision of the input range into buckets while fitting the input distribution using monotone cubic spline interpolation. The core challenge is implementing this method in a space-efficient manner while ensuring strong worst-case guarantees.
journalyear: 2025conference: ; ; copyright: none
1. Introduction
Data sketching has become one of the main tools for dealing with massive data volumes. Sketches provide a scalable way to extract key features from large datasets, enabling real-time analysis at streaming speed or massively parallel processing of distributed datasets, in applications such as network monitoring, machine learning, privacy, or bioinformatics.
Approximating order statistics is one of the central sketching problems. The goal is to process a massive dataset that is possibly distributed and summarize it in a small space into a sketch that is updatable with new observations, mergeable, and provides highly accurate estimates for the median, percentiles, and their generalization, quantiles. Equivalently, we can estimate data distribution by approximating ranks, i.e., the number of input items no larger than a given element $y$. Quantile sketches have been applied, e.g., to monitoring latencies (Tene, 2015) or for detecting anomalies (Siffer et al., 2017).
Streaming quantile estimation has thus received a significant attention in research community. From the theory point of view, it has been solved to a large extent. Namely, optimal algorithms are known for the uniform error, such as the Greenwald-Khanna sketch (Greenwald and Khanna, 2001; Cormode and Veselý, 2020) and KLL (Karnin et al., 2016). However, most of these theory-based quantile summaries are rarely used in practice. Instead, practitioners often use algorithms that perform significantly better on real-world datasets but have no guarantees. Most notably, $t$-digest (Dunning and Ertl, 2019) is widely used and, according to (Dunning, 2021), it has been adopted by major tech companies (reportedly Microsoft, Facebook, Google) and open-source projects (Elasticsearch, TimescaleDB, Apache Dubbo, etc.). At the same time, no matter how large the $t$-digest is, its error can be arbitrarily bad (Cormode et al., 2021).
Here, we introduce a new quantile sketch, called SplineSketch, that not only gets the best of both worlds – theoretical guarantees and very high accuracy in practice – but also significantly outperforms the state of the art, including $t$-digest (Dunning and Ertl, 2019). It processes any numerical input and requires a single parameter $k$ that determines its space usage, which is $O(k)$ memory words for processing the input and $16\cdot k$ bytes when serialized for storage. The sketch deterministically guarantees a theoretically near-optimal error, close to the best possible error of $O(n/k)$ with $k$ memory words; namely, the error is $O(n/k\cdot\log\alpha)$, where $\alpha$ is the aspect ratio, defined as the ratio of the largest and the smallest difference between distinct items in the dataset (e.g., for fixed-point numbers, $\alpha$ is bounded by the universe size). At the same time, the accuracy of our sketch in practice is in fact typically 10–200 better than $n/k$. Furthermore, we prove our sketch is fully mergeable while retaining the guarantees, i.e., we analyze it in the most general setting when the sketch is constructed by an arbitrary sequence of pairwise merge operations.
(a) HEPMASS dataset (Whiteson, 2016)
(b) Power dataset (Hebrail and Berard, 2006)
(c) Books dataset (Kipf et al., 2019; Marcus et al., 2020)
Figure 1. Average error of SplineSketch, KLL, MomentSketch, and $t$-digest on three real-world datasets (Sec. 5).
[…]
Overall, we demonstrate that besides mergeability, worst-case guarantees, and high accuracy in practice, our approach offers flexibility, making it suitable for a range applications.
1.1. Overview of our approach
We develop a new technique for designing quantile sketches that draws inspiration from both theoretical and practical approaches. The core of our approach is in maintaining a subdivision of the input range into buckets in a way that better adapts to the input distribution. Namely, SplineSketch consists of thresholds $\tau_{1},\cdots,\tau_{k}$ and then $(\tau_{i-1},\tau_{i}]$ is the $i$-th bucket, with $\tau_{0}=-\infty$. For each bucket, we store a counter $b_{i}$ that represents an approximate number of items that lie in $(\tau_{i-1},\tau_{i}]$. Furthermore, we use cubic spline interpolation in order to answer queries more accurately in practice.
Figure 2. SplineSketch illustration, with the values and thresholds $\tau_{i}$ on the $x$ axis and the rank space $[0,n]$ on the $y$ axis. The true cumulative distribution function of the data is depicted as a red dashed curve. The green dots correspond to the prefix sums of bucket counters, i.e., the counter $b_{i}$ equals the difference of the $y$ values of the the green dots at $\tau_{i}$ and $\tau_{i-1}$. The solid black curve is an interpolation over the green dots.
Answering queries using cubic splines.
The estimated rank of any threshold $\tau_{i}$ is $\sum_{j=1}^{i}b_{i}$. It is however not clear how to answer rank queries that are not at one of the bucket boundaries $\tau_{i}$. Most of the previous work would just instead query the closest value $\tau_{j}$ and return that as the answer, or perform a linear interpolation. However, we seek to do better than that. To this end, we use cubic interpolation. Specifically, if we are querying $x$, such that $\tau_{i}<x<\tau_{i+1}$, we estimate the ranks $\hat{r}_{i-1},\cdots,\hat{r}_{i+2}$ of $\tau_{i-1},\cdots,\tau_{i+2}$. We then use a piecewise cubic Hermite interpolating polynomial (PCHIP) in order to interpolate between $\hat{r}_{i}$ and $\hat{r}_{i+1}$. Namely, we evaluate this interpolating polynomial at $x$ and return it as our estimate.
Maintaining the buckets.
The most important choice we need to make is what the buckets should be. In general, we will set $\tau_{0}$ and $\tau_{k}$ to be the smallest and the largest item, respectively, that we have seen so far. The main part is adjusting the buckets by splitting some buckets and joining pairs of adjacent buckets. When we join two adjacent buckets, we set the new counter to the sum of the two buckets’ counters. In the split procedure, we split the bucket in the middle. We then need to choose how to split the count into the two new sub-buckets. To this end, we perform a query at the midpoint (as described above) and we divide the count so as to make the value at the midpoint as returned by the query. It remains to specify how we choose which buckets to join and split. When describing this here, we ignore the question of efficiency. In the end, we will have a buffer and perform updates in batches, allowing for low amortized update time.
As an invariant, to ensure we always have $k$ buckets, whenever we split a bucket, we join two buckets, which we choose as described below. There are two rules to choose a bucket to split: First, throughout the execution, we ensure that all buckets have counter $O(n/k)$, and we split a bucket when it gets too big. We use these bucket size bounds, together with other invariants, to show the worst-case error bounds. Intuitively speaking, we show that no item can be part of many splits and that each split can only contribute to the error by at most the size of the bucket the item belongs to.
Second, while we want to get good error bounds in the worst case, we seek to perform significantly better in practice. To this end, we also introduce for each bucket a number which we call heuristic error, which is our guess for the actual error of the cubic spline inside the bucket. The underlying reason is that how the data distribution function changes is captured by the derivatives of the distribution. For example, suppose that the data items are i.i.d. from an unknown distribution. If the pdf $f$ does not change too much around a bucket, i.e., the pdf’s derivative $f^{\prime}$ is small, the error should be low as well. Of course, we do not know $f$ and data may not be i.i.d., but an estimate of the pdf’s derivative around a bucket serves as a heuristic to the true error. This number is not stored but we compute it when we need it. If the maximum heuristic error can be decreased substantially by splitting one bucket and joining two other adjacent buckets, then we perform the split and join of these buckets. The exact function that we use as the heuristic error is discussed in Section 3.1 (see Equantion 2).
When we need to perform a join (typically when we split a bucket), we would like to join two adjacent buckets whose joining would result in a bucket with the lowest possible value of heuristic error. However, this may not be possible for two reasons. First, it may violate the bucket size bound – this can be easily fixed by only considering those pairs of buckets where the bucket size bound will not be violated after the join. In fact, we do not join two buckets if the resulting bucket would be, say, 75% full. Second, in some situations this may lead to too many joins and splits on a single bucket, which would invalidate our worst-case bounds. To this end, we divide the input into epochs, where one epoch is defined as time during which the input size increased by a fixed constant factor. We then keep a bitmap to ensure that a bucket that was created by a split will not be joined during the same epoch, that is, we protect thresholds of newly created buckets from joining for a certain time period. This allows us to prove the worst-case bounds.
Merging two sketches.
The merge operation of our sketch is conceptually straightforward: First, we estimate the ranks of thresholds of each sketch in the other sketch. Then we take the union of the two sets of thresholds, disregarding duplicate thresholds or thresholds that are too close to another threshold, which results in a sketch with possibly too many buckets. We set the counters of new buckets based on their estimated ranks in both sketches. Finally, we join enough pairs of adjacent buckets based on their heuristic error so that the resulting sketch has the desired number of buckets. To ensure worst-case guarantees, the protection of thresholds from joining is inherited from the larger sketch if merging two sketches of substantially different size.
Resizing the sketch.
Using joining or splitting the buckets allows us to easily change the number $k$ of buckets. Namely, when we want to make the sketch smaller, we just perform the appropriate number of buckets joins without any splits, whereas for increasing the sketch size, we perform splits without joins. This demonstrates the flexibility of our approach, and is useful in practice, e.g., when we allocate more memory for processing the input and then compress the sketch into much smaller size.
1.2. Related work
Quantile summaries are intensively investigated since the 90s, with the pioneering work of Munro and Paterson from 1978 (Munro and Paterson, 1980) who showed that exact quantile selection in sublinear space is impossible. In the comparison-based model, Greenwald and Khanna (Greenwald and Khanna, 2001) designed a deterministic algorithm that stores $O(\varepsilon^{-1}\cdot\log\varepsilon n)$ stream items and guarantees ±$\pm\varepsilon\cdot n$ uniform error; this bound is optimal among deterministic comparison-based quantile summaries (Cormode and Veselý, 2020). However, the GK summary is not known to be mergeable while retaining its space bounds (Agarwal et al., 2013), despite that the intricate analysis from (Greenwald and Khanna, 2001) was recently simplified and generalized to weighted updates (Assadi et al., 2023; Gribelyuk et al., 2024). Randomization allows to remove the logarithmic dependency on the stream length $n$, and after a sequence of improvements (Manku et al., 1999; Luo et al., 2016; Felber and Ostrovsky, 2015), Karnin, Lang, and Liberty (Karnin et al., 2016) developed an optimal randomized algorithm that stores just $O(\varepsilon^{-1})$ items (with a constant probability of a too large error). Beyond the comparison-based model, $q$-digest (Shrivastava et al., 2004) builds an implicit binary tree over the items’ universe $\mathcal{U}$ and stores a subset of the tree nodes. It requires space of $O(\varepsilon^{-1}\cdot\log|\mathcal{U}|)$ and a foreknowledge of $\mathcal{U}$. Very recently, the space usage of $q$-digest was decreased to the optimal $O(1/\varepsilon)$ memory words, by packing more nodes and their associated counters into one memory word (Gupta et al., 2024).
Another line of work focused on the stronger relative (multiplicative) error guarantee, which requires that the error for an item of rank $r$ is at most ±$\pm\varepsilon\cdot r$. For deterministic comparison-based algorithms, there is a gap of $O(\log\varepsilon n)$ between the merge-and-prune algorithm using space $O(\varepsilon^{-1}\cdot\log^{3}\varepsilon n)$ (Zhang and Wang, 2007) and the lower bound of $\Omega(\varepsilon^{-1}\cdot\log^{2}\varepsilon n)$ (Cormode and Veselý, 2020). Efficient randomized sketches with relative error have only appeared recently, namely, ReqSketch with space $O(\varepsilon^{-1}\cdot\log^{1.5}\varepsilon n)$ (Cormode et al., 2023) and its “elastic“ version achieving space close the information-theoretic lower bound of $\Omega(\varepsilon^{-1}\cdot\log\varepsilon n)$ up to asymptotically smaller factors (Gribelyuk et al., 2025); however, mergeability of the latter sketch is open. There is also an extension of $q$-digest to the relative error that uses space $O(\varepsilon^{-1}\cdot\log\varepsilon n\cdot\log|\mathcal{U}|)$ (Cormode et al., 2006); it is not known to be optimal.
For dynamic streams (i.e., with deletions), one can only achieve the uniform error guarantee in sublinear space, and the best quantile summary supporting deletions is the Dyadic CountSketch (Luo et al., 2016).
Finally, there are more heuristic approaches that do not aim for any worst-case guarantees and work only for numerical data, similarly as our algorithm. Specifically, $t$-digest (Dunning and Ertl, 2019; Dunning, 2021) performs an online one-dimensional $k$-means clustering by averaging data items into a given number of centroids, and interpolates linearly between the centroids. A downside of this approach is that many of the items summarized by a centroid may be smaller or larger than an adjacent centroid mean, leading under- or over-estimation. That is, the centroids are only ordered by their means, while the represented items may be far from ordered. In fact, one can impose high overlaps of represented items, leading to almost arbitrarily large error on adversarial instances (Cormode et al., 2021). Linear interpolation techniques were combined with the KLL sketch to improve its accuracy (Schiefer et al., 2023). In contrast to rank error, DDSketch (Masson et al., 2019) and UDDSketch (Epicoco et al., 2020) provide the relative value error (i.e., an item close to the desired quantile in the item space) based on maintaining a suitable exponential histogram. Finally, MomentSketch (Gan et al., 2018; Mitchell et al., 2021) is possibly the most compact summary as it consists of $k$ moments and log-moments of data items, for a small $k$ such as $k=15$. Upon a query, the sketch constructs a distribution with the same moments and log-moments, using the maximum entropy principle and Chebyshev polynomials.
2. Preliminaries
Throughout the paper, $n$ is the number of items currently summarized by the sketch. For a fixed multiset $S$ of real numbers and for any $x\in\mathbb{R}$ let $\mathrm{rank}(x)=|\{y\in S:y\leq x\}|$, i.e., the number of input items smaller or equal to $x$. For any $q\in(0,1]$, we call $y:=\min\{x:\mathrm{rank}(x)\geq q\cdot n\}$ the $q$-quantile of $S$. This definition implies that a method for estimating ranks may be used for estimating quantiles with the same accuracy, by performing binary search. The aspect ratio $\alpha$ of an input $x_{1},\dots,x_{n}$ is defined as $\alpha=(\max_{i}x_{i}-\min_{i}x_{i})/\gamma$, where $\gamma=\min_{i\neq j:x_{i}\neq x_{j}}|x_{i}-x_{j}|$ is the smallest difference of distinct items on input (assuming at least two distinct items are present).
We use a monotone variant of the Piecewise Cubic Hermite Interpolating Polynomial (PCHIP) (Fritsch and Butland, 1984; Fritsch and Carlson, 1980), which is an interpolation method that constructs a continuous function by fitting cubic polynomials between data points. Unlike standard cubic spline interpolation, PCHIP adjusts the first derivatives at each data point to preserve the shape and monotonicity of the original data. For an introduction at a greater detail, including the formulas for the interpolation, see (Wong, 2020). Since the ranks of the input items are non-decreasing, preserving monotonicity is an important feature of the interpolation, rendering PCHIP suitable for our application.
3. Description of SplineSketch
We provide a description of our algorithm, omitting implementation details, explained in Sec. 4. In fact, there are many possible implementations of the sketch that still satisfy the worst-case guarantee. Our sketch works for any numerical input, i.e., the input items are integers or floating-point numbers of bounded precision. The theoretical analysis of the worst-case error is deferred to Appendix  B.
For simplicity, we first present the sketch assuming there are no items of high frequency, namely appearing at least $n/k$ times in the input after inserting any number $n$ of items into SplineSketch of size $k$. We show how to remove this assumption in Sec. 3.3.
Buckets.
Fix an integer parameter $k\geq 6$ (as required in the analysis). The sketch primarily consists of $k$ distinct thresholds $\tau_{1}<\tau_{2}<\cdots<\tau_{k}$ and $k$ counters $b_{1},\ldots,b_{k}$, where $b_{i}$ is our estimate of the number of items that lie in the interval $(\tau_{i-1},\tau_{i}]$; we define $\tau_{0}=-\infty$. We call these intervals buckets and refer to a bucket by the index of its right threshold, so $(\tau_{i-1},\tau_{i}]$ is called ”bucket $i$” or the $i$-th bucket. We also frequently refer to $b_{i}$ as the size of bucket $i$ and we define the length of bucket $i$ as $\ell_{i}:=\tau_{i}-\tau_{i-1}$. Throughout the execution of the algorithm $\tau_{1}$ and $\tau_{k}$ are always set to be the minimum and the maximum item in the stream, respectively. This makes the first bucket special as its counter only stores the frequency of $\tau_{1}$.
We also maintain a buffer of size $O(k)$ that we use for faster processing of incoming items, i.e., low average update time, and one auxiliary bit-array of length $k$ for “protected thresholds” (Sec. 3.1). After processing the whole input, the buffer is merged into the buckets and the auxiliary bit-array can be discarded. We maintain the two invariants: All buckets counters plus the number of items in the buffer sum up to $n$ and that no bucket is empty.
Insertion and initial buckets.
When a new element arrives, we first put it in the buffer. When the buffer gets full, we call the consolidate method (Sec. 3.1) that may change bucket thresholds and merges the buffer into buckets. The only exception is the first time when the buffer gets full. In that case, we simply sort it, pick every $n/k$-th item, thus selecting $k$ items, and set the counters $b_{i}$ to exactly the number of items in the corresponding buckets. Since no item has frequency at least $n/k$, these selected items are distinct.
Query.
For a query at a threshold $\tau_{i}$, we return the sum of the counters $\sum_{j=1}^{i}b_{i}$. For $x$ strictly inside the $i$-th bucket, i.e. $\tau_{i-1}<x<\tau_{i}$, we use the PCHIP interpolation to estimate the rank of $x$; note that this works no matter whether $x$ appeared on input or not. We remark that computing this interpolation requires only the knowledge of $\tau_{i-2},\cdots,\tau_{i+1}$ together with corresponding prefix sums of bucket counters. The interpolation does not extrapolate, i.e., for $x<\tau_{1}$, it returns 0, and for $x>\tau_{k}$, it returns $\sum_{j=1}^{k}b_{i}$. Since one has to find the right bucket by binary search, one query takes time $O(\log k)$ provided that the prefix sums are precomputed.
3.1. Consolidating buckets
The main procedure of the sketch is the consolidate method. The method first counts the new number of items in each bucket if we were to merge the buffer into the buckets. Then, it keeps changing the buckets based on the projected new values of the counters, possibly removing some of the thresholds and creating new ones. After those changes, the buffer gets merged into the new buckets, meaning we again count for each bucket the number of items currently in the buffer that fall into that bucket and update the counters accordingly. Finally, the buffer gets emptied. The crux of the consolidate method is how to modify the thresholds, which we describe in the remainder of this subsection.
Joining and splitting buckets.
We have two basic operations that change bucket thresholds, join and split. A join takes two neighboring buckets $(\tau_{i-1},\tau_{i}]$ and $(\tau_{i},\tau_{i+1}]$ and turns them into a single bucket $(\tau_{i-1},\tau_{i+1}]$ with its counter set to $b_{i}+b_{i+1}$. Note that after joining, buckets are (implicitly) renumbered. That is, we actually set new thresholds $\tau^{\prime}$ and counters $b^{\prime}$ as $\tau^{\prime}_{i}=\tau_{i+1}$, $b^{\prime}_{i}=b_{i}+b_{i+1}$, and for any $j>i$, $\tau^{\prime}_{j}=\tau_{j+1}$ and $b^{\prime}_{j}=b_{j+1}$. The buckets up to $\tau_{i-1}$ do not change.
When splitting bucket $(\tau_{i-1},\tau_{i}]$, we replace it by two buckets $(\tau_{i-1},\frac{\tau_{i-1}+\tau_{i}}{2}]$ and $(\frac{\tau_{i-1}+\tau_{i}}{2},\tau_{i}]$. The bucket counters are set as follows: We use the query method to get estimated ranks $\hat{r}$ at $\tau_{i-1},\frac{\tau_{i-1}+\tau_{i}}{2},\tau_{i}$. We then set the counters for the two new buckets to $\hat{r}(\frac{\tau_{i-1}+\tau_{i}}{2})-\hat{r}(\tau_{i-1})$ and $\hat{r}(\tau_{i})-\hat{r}(\frac{\tau_{i-1}+\tau_{i}}{2})$, respectively. Other buckets are again implicitly renumbered.
Observe that every split increases and every join decreases the number of buckets by one. Thus, we accompany every split with joining two buckets, which ensures there are always $k$ buckets (unless a change of $k$ is desired). We argue in Appendix B.2 that this can always be done in a way that preserves worst-case guarantees.
Handling new minima and maxima.
If the buffer contains a new minimum $\tau^{\prime}_{1}$ (an item smaller than the previous minimum $\tau_{1}$), we add a new bucket $(-\infty,\tau^{\prime}_{1}]$ before all of the existing buckets, set its counter to the frequency of $\tau^{\prime}_{1}$ in the buffer, and we also increase the counter of the bucket $(\tau^{\prime}_{1},\tau_{1}]$, where $\tau_{1}$ is the previous minimum, by the number of buffer items in this interval. A new maximum $\tau^{\prime}_{k}$ is handled similarly, by adding a new bucket $(\tau_{k},\tau^{\prime}_{k}]$ and setting its counter to the number of buffer items in this interval. For each of these at most two new buckets, we perform a join of two buckets. The rules for selecting which pairs buckets to join are described in the following paragraphs. Note that these at most two new buckets may be split during the consolidation that created them.
Maintaining bounds on bucket counters and heuristic error.
We maintain two invariants, the first of which is used to prove worst-case bounds while the second ensures good performance in practice.
In order to obtain low worst-case uniform error we need to ensure that no bucket has too large counter. To this end, we require that all bucket counters are bounded by $O(n/k)$, specifically
(1) $\forall i=1,\dots,k:\,b_{i}\leq C_{b}\cdot n/k$
for some constant $C_{b}>1$ that will be determined in the analysis in Section B.3.
Secondly, we define a heuristic error for each bucket, which is intended to capture “hard” parts of the input, where the input distribution changes a lot, namely, where the derivative of the empirical PDF is large. For such buckets, the PCHIP interpolation may have larger error than for buckets with low derivative of the PDF. Since the empirical PDF in the bucket $i$ is $b_{i}/\ell_{i}$, we use the following heuristic error for bucket $i$, which approximates the second derivative of the empirical CDF, normalized by length squared:
(2) $\max\left(\frac{\left|\frac{b_{i}}{\ell_{i}}-\frac{b_{i-1}}{\ell_{i-1}}\right|}{\ell_{i}+\ell_{i-1}}\,,\,\frac{\left|\frac{b_{i+1}}{\ell_{i+1}}-\frac{b_{i}}{\ell_{i}}\right|}{\ell_{i+1}+\ell_{i}}\right)\cdot\ell_{i}^{2}.$
For the edge cases $i=1$ and $i=k$, we make the following assumptions: $b_{0}=b_{k+1}=0$, $\ell_{0}=\ell_{1}$, and $\ell_{k+1}=\ell_{k}$; that is, we add “virtual” buckets 00 and $k+1$ with zero counters and a length equal to the length of the adjacent (real) bucket.
We now describe how we use the bucket bound (1) and heuristic error (2) when performing the splits and joins. Each time we split a bucket $i$, we choose a pair of adjacent buckets to join in order to keep the total number of buckets the same. This pair is selected so that after joining it would yield a bucket with the lowest heuristic error, of all joinable pairs, defined as follows, which are not overlapping bucket $i$.
Definition 1.
We say that a pair of adjacent buckets is joinable if both of the following conditions hold:
  1. (i)
    the bucket resulting from the join would not violate the bucket size bound (1) or be relatively close to it, specifically it would have counter at most $0.75\cdot C_{b}\cdot n/k$, and
  2. (ii)
    any of the two buckets were not subject to a split recently; specifically, threshold $\tau_{i}$ is not protected as explained below.
We perform splits as follows. First, we take all buckets that exceed the size bound (1). We split them one by one, always performing a join satisfying conditions of Definition 1 for each split; in Lemma 2 we show that at least one pair of adjacent buckets can be joined. We also join at most two further pairs of joinable buckets if the buffer contains a new minimum or maximum, each creating a new bucket. Note that a bucket may need to be split repeatedly during one call of the consolidate method, namely, if it exceeds the bucket bound (1) substantially.
After all these necessary splits and joins, we consider a bucket $i$ with a high heuristic error (2), and we split it if its error is more than $\gamma$ times greater than the lowest heuristic error of a bucket resulting from joining a joinable pair, for a parameter $\gamma>1$ (say, $\gamma=1.5$). Furthermore, we do not perform a split of a bucket $i$ due to the heuristic error if there are less than $k/3+2$ pairs of adjacent buckets that can be joined, which is needed to prove that there is a joinable pair of buckets; see Lemma 2. Finally, if the counter of bucket $i$ is substantially below the bound, say, $b_{i}\leq(C_{b}/100)\cdot n/k$, we also do not split bucket $i$, thus avoiding to split almost empty buckets. We summarize the conditions for splitting:
Definition 2.
We say that a bucket $i$ is splittable if one of the following conditions holds:
  1. (i)
    bucket $i$ violates the bucket bound (1), i.e., $b_{i}>C_{b}\cdot n/k$, or
  2. (ii)
    there are at least $k/3+2$ joinable pairs of buckets, $b_{i}>(C_{b}/100)\cdot n/k$, and the heuristic error (2) of bucket $i$ is $\gamma>1$ times larger than the heuristic error of a joinable pair of buckets (after joining) not overlapping bucket $i$.
We repeat finding a splittable bucket $i$ and a joinable pair of buckets not overlapping bucket $i$, performing the split and join as long as there are splittable buckets. Lemma 2 shows that if there is a splittable bucket $i$, then there is a joinable pair of buckets not overlapping bucket $i$.
Protected thresholds and epochs.
The above description yields a sketch with high accuracy in practice, but contrived instance may cause a large error, much larger than $n/k$. Basically, the issue is that one bucket may be split and joined repeatedly, while accumulating error; in the extreme case, one bucket may undergo $\Omega(k)$ splits and joins, and each time, the error at the threshold increases by $\Omega(n/k)$, leading to $\Omega(n)$ error. With the small overhead of $k$ bits when processing the input, we show how to avoid such adversarial repeated splits and joins at a single location by protecting thresholds of buckets that were recently created by a split.
Specifically, we divide the input into epochs. The first epoch has length $\Theta(k)$ and each successive epoch starting at time $t$ has length $0.25\cdot t$. When we split a bucket $(\tau_{i-1},\tau_{i}]$, we mark all three thresholds $\tau_{i-1},\tau^{\prime}_{i},\tau_{i}$ as protected, where $\tau^{\prime}_{i}=(\tau_{i-1}+\tau_{i})/2$ is the new threshold. We store a bit vector of length $k$, with the $i$-th bit set to 1 when the $i$-th threshold is protected. When a new epoch starts, we reset all thresholds to the default unprotected state by zeroing the bit vector. As specified in Definition 1, we never join buckets that have a protected threshold between them.
Note that we may still split a single bucket multiple times during one epoch, or even during one call of the consolidate method (namely, if the bucket receives a large number of items). Similarly, a single bucket may be joined multiple times during one consolidation, but a combination of splitting and joining is not allowed – bucket which was created by a split cannot be joined with another bucket in the same epoch.
Resizing sketch during consolidate.
In fact, the consolidate method allows to easily change the number $k$ of buckets, thus making the sketch larger or smaller. Changing $k$ to $k^{\prime}$ buckets is done in a straightforward way: If we increase the sketch size, i.e., $k^{\prime}>k$, we perform $k^{\prime}-k$ splits without any join, first dealing with buckets with exceeded bound and then in the order of the heuristic error. On the other hand, if $k^{\prime}<k$, after performing the necessary splits due to exceeding the bound, we execute $k-k^{\prime}$ joins in the order of heuristic error, on bucket pairs satisfying Definition 1 and without performing any splits. These separate splits or joins may be followed by performing consolidation in a normal way, where we always join after splitting and vice versa. If $k^{\prime}$ is substantially different from $k$, say $|k^{\prime}-k|>0.25\cdot k$, we also end the current epoch, i.e., reset the protection bit vector. Note that our analysis in Appendix B does not account for sketch resizing.
3.2. Merging Two SplineSketches
We merge two sketches $S_{1}$ and $S_{2}$ with $k_{1}$ and $k_{2}$ buckets, respectively, as follows. Without loss of generality, at least as many items have been inserted into $S_{1}$ as into $S_{2}$. First, we create a buffer that contains all the items from the buffers of the two sketches we are merging.
Second, we take the union of all the thresholds $\tau$ from both sketches. After removing duplicates, this gives up to $k_{1}+k_{2}$ thresholds. For each of these thresholds $\tau$, we query for its rank in both sketches and add these two ranks to get the rank of $\tau$ in the merged sketch. We set the counter of each bucket to be the difference between the ranks of the endpoints.
Third, the new sketch inherits the bit vector for threshold protection from $S_{1}$ as it summarizes a larger input than $S_{2}$. If the total input size of the new sketch is larger than the epoch end for $S_{1}$, we set all thresholds to the unprotected state, starting a new epoch.
Finally, we run the consolidate method that merges a potentially overflown buffer to buckets and more importantly, reduces the number of buckets to $k_{1}$. Namely, we repeatedly take a joinable pair of adjacent buckets whose joining would result in a bucket with the lowest heuristic error and join it without any split before we have $k_{1}$ buckets.
3.3. Dealing with High-Frequency Items
The algorithm as described above works and satisfies the worst-case bound as long as all items’ frequencies are below $n/k$.
There are two ways to deal with high-frequency items. First, one can satisfy the assumption of no high-frequency items on any input by running the Misra-Gries (MG) sketch (Misra and Gries, 1982) of size $O(k)$ in parallel. That is, we first insert an incoming item $x$ into the MG sketch. If $x$ is stored by MG, then we do not add $x$ to SplineSketch; otherwise, we add $x$ to SplineSketch. The latter case may result in some items being ousted from the MG sketch, which we add to SplineSketch together with their frequencies in MG.
However, using MG or another heavy-hitter algorithm would make our sketch more complicated and slower. In the actual implementation, we do not maintain a heavy-hitter sketch in parallel, but rather try to recognize heuristically whether or not a bucket represents a single item of high frequency as it has a very small relative length. See Section 4.2 for details.
4. Implementation Details
Quantile queries.
So far, we focused on answering rank queries. In order to implement a quantile query, i.e., returning an estimated $r$-th smallest number for a given $r$, we invert the interpolation, and query the inverse function at $r$; the inverse exists since buckets are non-empty and the PCHIP interpolation is thus strictly increasing from $\tau_{1}$ to $\tau_{k}$. The inverse function can be numerically evaluated by using, for example, the Newton’s method or even just binary search. Note that our algorithm, similarly as q-digest, may return a number that did not appear in the stream, unlike comparison-based methods such as the KLL sketch (Karnin et al., 2016).
4.1. Implementations of the consolidate method
We describe two particular implementations of the consolidate method. The first achieves a low amortized update time, while the second is easier to implement and works well in practice.
Heap-based consolidate.
To obtain asymptotically fast consolidate method, one can use a combination of doubly-linked lists for maintaining buckets and two heaps, sorted by the bucket size and heuristic error, respectively. This approach is formalized in the following lemma. Since the buffer size is $O(k)$, we get that the amortized complexity of an update with heap-based consolidate is $O(\log k)$.
{restatable}
lemmaheapBasedConsolidate Consolidation can be implemented in $O(k\log k)$ time.
Proof.
We maintain buckets in a doubly linked list $L$. We use a queue $Q_{S}$ to keep track of buckets that need to be split due to violating bucket bound. We also use two heaps, $H_{S}$ and $H_{J}$, to keep track of buckets that might get split due to heuristic error and joinable pairs of buckets, respectively. $H_{S}$ is a max-heap with heuristic errors as keys, and $H_{J}$ is a min-heap with the keys being the heuristic errors after the join. In both heaps, every node contains a pointer to the corresponding bucket (or pair of buckets) in $L$. Conversely, every bucket in $L$ contains a (constant size) list of pointers to nodes of $H_{S}$ and $H_{J}$ containing that bucket. Both heaps can be initialized at the beginning of consolidate in $O(k)$ time.
After every split and join performed by consolidate method, we update the $Q_{S},H_{S}$, and $H_{J}$ according to the new bucket sizes. Since every split and join affect only the sizes of constant number of adjacent buckets, the update can be done in $O(1)$ time using the pointers that we maintain. The modifications of the keys and deletions of arbitrary nodes in the heaps can be done in $O(\log k)$ time using standard decrease/increase key operation.
Every split introduces a constant number of new protected thresholds and every join is accompanied with a corresponding split. Since the number of possible protected thresholds is bounded by $O(k)$, the total number of splits and joins performed by consolidate is also $O(k)$. Therefore, the total time complexity of the heap based implementation of consolidate is $O(k\log k)$.
Iteration-based consolidate.
In our implementation, we use practically more efficient consolidate method without the need for linked lists and heaps. We process the buckets in iterations such that each iteration splits or joins one bucket at most once and not both at the same time; note that more iterations are possibly needed as one bucket may be split or joined mutlple times during consolidation. We stop when the last iteration does not change any threshold. Throughout this process, we keep the buffer and the old buckets for more accurate setting of counters for the new buckets.
In each iteration, we first collect the buckets with size exceeding the bound (1) into a set $E$. We also keep a list of splittable buckets (Def. 2), sorted non-increasingly by their heuristic error, and a list of joinable pairs of buckets (Def. 1), sorted non-decreasingly by their heuristic errors after potentially performing their join. We select $|E|$ joinable pairs according to their order that do not overlap buckets in $E$ or a previously selected joinable pairs. Next, from the remaining buckets (not yet split or joined), we consider the splittable bucket with the largest heuristic error $h_{\mathrm{split}}$ (if any) and the joinable pair with the lowest heuristic error $h_{\mathrm{join}}$ and if $h_{\mathrm{split}}>1.5\cdot h_{\mathrm{join}}$, then we perform the split and join of these buckets. After collecting non-overlapping sets of buckets to split and join, we perform these splits and joins in one pass over the buckets, using the buffer and interpolation over the buckets before the consolidation; the usage of buckets before consolidation is important otherwise error may accumulate.
Finally, we remark that the multiplicative constant $C_{b}$ in the bucket bound (1) resulting from the analysis in App. B is too large for practical usage, and we just use $C_{b}=3$, increasing $C_{b}$ during an epoch if needed (i.e., if there are no joinable pairs when a bucket must be split due to exceeding the bound) and resetting it to $C_{b}=3$ at the end of each epoch; we note that such increases of $C_{b}$ happen rarely in practice.
4.2. Handling Frequent Items without a Heavy-Hitter Sketch
The potential presence of high-frequency items makes the implementation more convoluted. We chose not to use a separate heavy-hitter sketch to deal with frequent items as it would make SplineSketch slower when processing the stream while item frequencies are relatively small in many applications. Instead, we set a lower limit on the relative length of a bucket, depending on numerical precision $\delta$ of data. That is, a bucket $(\tau_{i},\tau_{i+1}]$ must have length $\tau_{i+1}-\tau_{i}\geq\delta^{\prime}\cdot\max\{|\tau_{i}|,|\tau_{i+1}|,\varepsilon\}$, where $\delta^{\prime}>\delta$ is a parameter and $\varepsilon>0$ is the smallest non-zero absolute value of an input item ($\varepsilon$ is intended to avoid too small buckets around 0). In our prototype implementations in Python and in Java with the double data type, we set $\delta^{\prime}=10^{-8}$. We do not split a bucket that would violate this relative length bound, even if it does not satisfy the size bound (1); this way, we avoid too many splits of a bucket with a frequent item.
We also deal with repeated thresholds when creating initial buckets by equally spaced selection (after the buffer gets full for the first time). Namely, we remove duplicate thresholds and instead, add new thresholds relatively close to the previously repeated thresholds (but not too close to violate the relative length bound) so that we end up with $k$ thresholds in total. This makes our approach robust to frequent items that appear from the beginning: The new thresholds will in fact heuristically guarantee that the frequent item gets its own bucket if it is frequent from the very beginning.
5. Experimental Evaluation
We have implemented SplineSketch as a prototype in Python and Java. Here, we provide a comparison with state-of-the-art quantile sketches for the uniform rank error with an available implementation, namely, $t$-digest (Dunning and Ertl, 2019), MomentSketch (Gan et al., 2018), and KLL (Karnin et al., 2016); see Appendix C for details. We run the sketches on three real-world and ten synthetically generated datasets (App. C). We measure the average and maximum absolute rank errors for $10^{5}$ evenly spaced queries, together with average update and query times. The sketch size is measured in bytes when serialized on disk for storage without supporting data structures such as buffers.
(a) Uniform distribution
(b) Pareto distribution
(c) Gumbel distribution
(d) Loguniform distribution
(e) Normal distribution with a large parameter change after $n/2$ items
(f) Normal distribution with $n/2$ items, followed by $n/2$ high-frequency items
Figure 3. Average rank error (log-scale) on synthetic datasets with $n=10^{8}$ items depending on the sketch size in bytes. The average is taken over $10^{5}$ evenly spaced queries.
Accuracy-space trade-off.
Figs. 1 and 3 show the average errors of the sketches depending on their size in bytes; see App. D for additional plots with average error (Fig. 8) and plots with the maximum error (Fig. 9). In all these plots, the error is presented in absolute terms, i.e., it is the absolute value of the difference between the true rank and the estimated rank of a query item, averaged or maximized over $10^{5}$ queries.
Overall, SplineSketch consistently provides the best accuracy if given sufficient size, namely $k\geq 100$, even though in rare cases or for small sketch sizes, $t$-digest or MomentSketch are better. In more detail, compared to $t$-digest, SplineSketch typically achieves 2-20 times smaller error, and in some cases, such as for normally distributed data, up to 100 times. On the other hand, there were cases in which $t$-digest performed better, such as for normal distribution followed by high-frequency items or in general, for datasets with skewed frequency distributions. In several other cases, their error was similar, e.g., the maximum error for signed loguniform distribution with large exponents. We note that the final $t$-digest size is somewhat unpredictable as the final number of centroids does not match the compression parameter given to it, sometimes by 66%. This demonstrates that $t$-digest does not use the space budget efficiently, unlike SplineSketch which always has $k$ buckets.
MomentSketch is mainly intended for use with very small size, typically up to $k=15$ moments and log-moments, when it often gives the best accuracy by 1-2 orders of magnitude. It works the best on many smooth distributions (uniform, normal, loguniform, etc.). However, its performace is significantly affected by non-smoothness as witnessed on the real-world datasets. Moreover, one cannot increase accuracy by increasing $k$ because of numeric issues. Furthermore, MomentSketch query procedure failed with an exception on many datasets, specifically on signed loguniform distribution with large exponents for any $k$ and on other datasets for large $k$. Finally, the KLL sketch is worse than SplineSketch typically by 1-3 orders of magnitude as it uses no interpolation (cf. (Schiefer et al., 2023) for KLL with interpolation; however, no implementation is available). Nevertheless, as KLL is comparison-based, KLL’s error depends just on the data size, not on their particular distribution, and decreases linearly with KLL’s size.
Figure 4. Log-log plot with time per update in $\mu$s depending on data size with fixed sketch size on normally distributed data.
Figure 5. Log-log plot with time per query in $\mu$s depending on the number of queries on $n=10^{8}$ samples from normal distribution.
Update and query times.
[…]
6. Conclusions and Discussion
In this work, we designed a new deterministic quantile sketch which has fixed memory consumption $O(k)$. The main technical contribution is in careful maintenance of bucket thresholds and counters in order to obtain extremely accurate rank and quantile estimates, with uniformly bounded worst-case error, achieving a near-optimal accuracy-space trade-off. We have proven that our sketch is fully mergeable, rendering it suitable for parallel or distributed processing of massive datasets.
Besides very high accuracy on various datasets, the key advantage of our approach is flexibility. For instance, it is possible to resize the sketch, making it smaller or larger depending on how the memory resources change (e.g., one can use larger size $k$ for input processing than for storage). More importantly, if the user has a prior knowledge about the data distribution, it is possible to preset the initial thresholds of SplineSketch based on this prior distribution and aim for even lower error. Furthermore, if the prior knowledge turns out to be wrong, our sketch adapts to the new distribution, similarly as we have shown in Fig. 3(e).
One of the challenges when implementing SplineSketch is dealing with high-frequency items. We have tweaked the implementation to deal with them if they arrive at the beginning of processing a stream; however, if a high-frequency item starts occurring in the middle, it is unclear how to maintain the thresholds and counters in a right way. Still, if one processes inputs with such items, it is possible to deal with them by using a standard heavy-hitter algorithm, such as the MG sketch (Misra and Gries, 1982). Our implementation of SplineSketch is so far a prototype (in Python and Java) intended to demonstrate the advantages of our approach. Due to the intricacies of maintaining the buckets, our sketch is slower in processing the input stream than KLL or MomentSketch; however, a well-optimized Java version may be significantly faster than $t$-digest.
From the theoretical point of view, we ask whether one can modify our algorithm so that the error bounds have better dependency on $\alpha$. We note that there is some evidence from the comparison-based lower bounds in (Cormode and Veselý, 2020) that a factor of $\log\log\alpha$ may be necessary without packing more buckets into one memory word, similarly as done for q-digest in (Gupta et al., 2024). The accuracy in practice, despite being very high already, may also be improved, and it is not clear where is the accuracy limit of quantile summaries on a real-world datasets.
Our work opens up a new direction for future work in quantile streaming estimation. In particular, we believe that our techniques have a potential to yield much better quantile sketches with relative error guarantees, capturing the distribution tails more accurately than the median. However, achieving the relative error is substantially harder as witnessed by several major open problems in the theory of relative-error quantile sketches (cf. (Cormode and Veselý, 2020; Gribelyuk et al., 2025)).
Acknowledgments.
A. Łukasiewicz and P. Veselý were partially supported by the ERC CZ project LL2406 of the Ministry of Education of Czech Republic. Part of this work was done while A. Łukasiewicz was a PhD student at the University of Wrocław. J. Tětek was supported by the VILLUM Foundation grant 16582. This research was partially funded from the Ministry of Education and Science of Bulgaria (support for INSAIT, part of the Bulgarian National Roadmap for Research Infrastructure). Part of this work was done when J. Tětek was visiting at the University of Wrocław and at Charles University. Part of this work was done while J. Tětek was employed at BARC, University of Copenhagen. P. Veselý was partially supported by Czech Science Foundation project 24-10306S and by Center for Foundations of Modern Computer Science (Charles Univ. project UNCE 24/SCI/008).
Appendix A Bounds on Middle Point of Monotonic Cubic Polynomials
The following lemma provides a key technical property of PCHIP interpolation that we use subsequently in the analysis of our algorithm.
Lemma 1.
Let $P(x)$ be a cubic polynomial such that $P(0)=0$, $P(1)=1$, and $P(x)$ is non-decreasing on the interval $[0,1]$. Then, $P\left(\dfrac{1}{2}\right)\in[1-\beta,\beta]\approx[0.066,0.933],$ where $\beta=\frac{1}{2}+\frac{\sqrt{3}}{4}\approx 0.933$.
Proof.
We prove that $P(\dfrac{1}{2})\leq\frac{1}{2}+\frac{\sqrt{3}}{4}$. The other inequality then holds by symmetry, that is, considering the polynomial $Q(x)=1-P(1-x)$ which is also non-decreasing and $Q(\dfrac{1}{2})=1-P(\dfrac{1}{2})$. From now on, we focus on proving that $P(\dfrac{1}{2})\leq\frac{1}{2}+\frac{\sqrt{3}}{4}$.
Let us denote the coefficients of $P$ by $a,b,c$ and $d$, that is,
$P(\lambda)=a\lambda^{3}+b\lambda^{2}+c\lambda+d\,.$
Because $P(0)=0$, we have that $d=0$. At the same time, because $P(1)=1$, we have that $a+b+c=1$. From the assumption that the polynomial is non-decreasing, we also have that its derivative is non-negative, meaning that for any $\lambda\in[0,1]$,
$3a\lambda^{2}+2b\lambda+c\geq 0\,.$
At $1/2$, the polynomial evaluates to
$f(a,b,c)=a/8+b/4+c/2\,.$
We need to prove that under the constraints we mentioned, it holds that
$f(a,b,c)\leq\frac{1}{2}+\frac{\sqrt{3}}{4}$
First, using the constraint $a+b+c=1$, we express $c$ as $c=1-a-b$. Substituting this into the inequality $3a\lambda^{2}+2b\lambda+c\geq 0$, we obtain $3a\lambda^{2}+2b\lambda+(1-a-b)\geq 0$. Simplifying this expression gives:
(3) $(3\lambda^{2}-1)a+(2\lambda-1)b+1\geq 0\,.$
Next, we choose $\lambda=\dfrac{3+\sqrt{3}}{6}$, which lies within $[0,1]$. We compute the coefficients:
$3\lambda^{2}-1=\dfrac{\sqrt{3}}{2},\quad 2\lambda-1=\dfrac{\sqrt{3}}{3}\,.$
Substituting these values into inequality (3) gives:
$\dfrac{\sqrt{3}}{2}a+\dfrac{\sqrt{3}}{3}b+1\geq 0\,,$
which simplifies to:
(4) $\sqrt{3}\left(\dfrac{a}{2}+\dfrac{b}{3}\right)+1\geq 0\,.$
We now express $f(a,b,c)$ in terms of $a$ and $b$. Since $c=1-a-b$, we have:
(5) $f(a,b,c)=\dfrac{a}{8}+\dfrac{b}{4}+\dfrac{1-a-b}{2}=\dfrac{1}{2}-\dfrac{3a}{8}-\dfrac{b}{4}\,.$
To relate inequality (4) to $f(a,b,c)$, observe that inequality (4) implies:
$\dfrac{a}{2}+\dfrac{b}{3}\geq-\dfrac{1}{\sqrt{3}}.$
Multiplying both sides by $-\dfrac{3}{4}$ given:
$-\left(\dfrac{3a}{8}+\dfrac{b}{4}\right)\leq\dfrac{3}{4\sqrt{3}}.$
From Equation (5), we have:
$f(a,b,c)=\dfrac{1}{2}-\left(\dfrac{3a}{8}+\dfrac{b}{4}\right)\,.$
Using the inequality derived above, we obtain:
$f(a,b,c)\leq\dfrac{1}{2}+\dfrac{3}{4\sqrt{3}}=\dfrac{1}{2}+\dfrac{\sqrt{3}}{4}\,.$
By scaling on both x- and y-axes, Lemma 1 implies that using PCHIP interpolation (or any other monotone cubic interpolation for that matter) for splitting buckets in the middle of the gap between two thresholds gives a pair of two buckets with counters smaller than the original bucket counter by at least a constant factor. This technical property is further used in our analysis in Appendix B.
Appendix B Proof of Uniform Error Guarantees and Full Mergeability
We provide a formal analysis of SplineSketch, showing that it guarantees the uniform error (Theorem 3). We first prove that the algorithm can always find a pair of buckets to join when we need to split a bucket (Lemma 2), showing that the algorithm is well-defined. We show both in the most general mergeability setting, i.e., when the sketch is created by an arbitrary sequence of pairwise merge operations executed on single data items (Section B.1). We only assume that all sketches in the process have the same size $k$. The same properties then hold in the streaming setting, which can be simulated by always merging the sketch with a single-item sketch consisting of the buffer only.
In the subsequent analysis, we assume that $k\geq 6$. The precise constant factors in the analysis depend on the interpolation used. Let $\beta>0$ be such that after splitting a bucket, the counters of both new buckets are smaller than the original counter by at least a factor of $1-\beta$. By Lemma 1, we have that $\beta=\frac{1}{2}-\frac{\sqrt{3}}{4}\approx 0.066$ for any monotonic cubic interpolation used to answer queries in the algorithm. For the linear interpolation we have $\beta=0.5$.
B.1. Mergeability setting
For analyzing sketches created by a sequence of pairwise merge operations, it is convenient to define a binary tree $T_{0}$ with $n$ leaves, each representing one data item and inner nodes with two children corresponding to the merge operations. Additionally, we will use inner nodes representing the operations with buckets, namely splitting a bucket or joining a pair of adjacent buckets. That is, we define $T_{0}$ with root representing the final sketch, leaves corresponding the single data items, and inner nodes of three types: (1) merge-nodes with two children, representing merge operations such that the sketch represented by a merge-node is obtained by the union of buckets from the two source sketches represented by the children, as described in Section 3.2, and (2) split-nodes representing one split of a bucket, and (3) join-nodes representing one join of two adjacent buckets; split-nodes and join-nodes have a single child. Each node represents the sketch after the operation is performed. (A technicality is that the number of buckets may possibly be $k-1$ or $k+1$ after a join or split.) Other operations inside the sketch, namely creating initial buckets by equally spaced selection and merging buffer into buckets, need not be represented in $T_{0}$ for our analysis (while equally spaced selection affects the error, it is in total only by $n/k$).
We further label the inner nodes by epochs. Namely, let $n_{0}:=\Theta(k)$ be the initial buffer size and $n_{j}:=1.25\cdot n_{j-1}$ for $j>0$ be the epoch ends. We a call a node $t$ of $T_{0}$ a $j$-node if $j$ is the smallest integer $j\geq 0$ such that the number of items summarized by the sketch represented by $t$ is at most $n_{j}$. Clearly, the labels are non-decreasing on any leaf-to-root path. Moreover, merging two sketches represented by $j$-nodes for $j>0$ results in a sketch represented by a merge-node labeled by $j^{\prime}>j$ as the two source sketches summarize a similar number of items. We thus obtain:
Observation 1.
For any epoch $j>0$, the $j$-nodes of $T_{0}$ form a disjoint union of paths.
The 0-nodes form possibly more complicated subtrees. However, sketches represented by 0-nodes have no buckets and are exact as they store all their items in the buffer.
B.2. Existence of Joinable Pair of Buckets
In describing the consolidate subroutine (Section 3.1), we assumed that there is always a pair of buckets that we can join if a bucket counter exceeds the bound (1). Now we show that it holds.
Lemma 2.
Consider any SplineSketch represented by a node $t$ of the merge tree $T_{0}$, and suppose that there is a splittable bucket $i$ according to Definition 2. Then there exists a joinable pair of adjacent buckets, according to Definition 1.
Proof.
First, observe that if bucket $i$ is splittable but does not exceed the bucket bound (1), there is a joinable pair of adjacent buckets by Definition 2. In the following, we thus assume that bucket $i$ exceeds the bound.
Note that a pair of adjacent buckets is not joinable for one of two reasons: (1) due to bucket bounds, namely that the resulting bucket would have counter exceeding $0.75\cdot C_{b}\cdot n/k$ or (2) due to protected thresholds. Let $n_{t}$ be the number of leaves in the subtree of $t$, i.e., the number of data items summarized by the sketch in $t$. Observe that there are at most $62n_{t}/(0.75\cdot C_{b}\cdot n_{t}/k)=\frac{8}{3}k/C_{b}\leq k/$ 6 pairs of buckets satisfying case (1), where we use $C_{b}\geq 16$. Thus, the main part of the proof is bounding the number of protected thresholds.
Let $j$ be the label of node $t$. By 1, $j$-nodes in the subtree of $t$ form a path $P_{j}$. If the sketches represented by nodes on $P_{j}$ all have at least $k/3+2$ joinable pairs of buckets, then the lemma holds. Consider the highest node $t^{\prime}$ on the path $P_{j}$ such that the sketch represented by $t^{\prime}$ has at least $k/3+2$ joinable pairs of buckets; we have that $t^{\prime}\neq t$, i.e., $t^{\prime}$ is below $t$, otherwise the lemma holds. Note that the sketch of the lowest node $u$ on $P_{j}$ has at most $k-1-k/6$ pairs that cannot be joined as there are no protected buckets, thus $t^{\prime}$ is well-defined since $k-1-k/6\geq k/3+2$ by $k\geq 6$.
Let $P^{\prime}_{j}$ be the subpath of $P_{j}$ consisting of nodes above $t^{\prime}$. By the definition of $t^{\prime}$, all of them have less than $k/3+2$ joinable pairs of buckets. Thus, the algorithm does not perform splits on sketches on $P^{\prime}_{j}$ due to the heuristic error, that is, it only splits due to exceeding the bucket bound.
To bound the number of split-nodes on $P^{\prime}_{j}$, we consider the following potential of each sketch represented by a node $v$:
(6) $\Psi(v):=\sum_{i=1}^{k}\max\left\{0,b_{i}-0.75\cdot C_{b}\cdot\frac{n_{v}}{k}\right\}\,,$
where $b_{i}$ is the counter of bucket $(\tau_{i-1},\tau_{i}]$ of the sketch of $v$. Clearly, $\Psi(v)\in[0,n_{v}]$ and moreover, the total increase of the potential on path $P^{\prime}_{j}$ is at most $n_{t}$. Next, observe that the potential does not increase by joining buckets (by Def. 1) and that every split at node $v$ of a bucket $i$ with counter $b_{i}>C_{b}\cdot n_{v}/k$ results in two buckets with counters at most $(1-\beta)\cdot b_{i}$. Since $\beta<0.25$, a split thus decreases the potential by at least
$\beta\cdot b_{i}>\beta\cdot C_{b}\cdot\frac{n_{v}}{k}\geq\beta\cdot C_{b}\cdot\frac{n_{t^{\prime}}}{k}\geq\beta\cdot C_{b}\cdot\frac{n_{t}}{1.25\cdot k}$
as every node $v$ on path $P^{\prime}_{j}$ has $t^{\prime}$ in its subtree and $n_{t}\leq 1.25\cdot n_{t^{\prime}}$ by the definition of node labels. Hence, the number of split-nodes on path $P^{\prime}_{j}$ is at most
$\frac{n_{t}}{\beta\cdot C_{b}\cdot\frac{n_{t}}{1.25\cdot k}}=\frac{1.25\cdot k}{\beta\cdot C_{b}}\,.$
We choose $C_{b}$ so that the number of split-nodes on path $P^{\prime}_{j}$ is at most $k/18$. As every split results in three protected thresholds, the number of thresholds that get protected on path $P^{\prime}_{j}$ is at most $k/6$. Since there are at least $k/3+2$ joinable pairs of buckets in node $t^{\prime}$, at most $k/6$ of these pair get non-joinable due to falling in case (1), and at most $k/6$ of them get non-joinable due to protected thresholds, there will be a joinable pair of buckets in node $t$. ∎
B.3. Proof of the Uniform Error Guarantee
We now formally prove the worst-case guarantee for SplineSketch, showing that it is fully mergeable [Agarwal et al., 2013, Cormode et al., 2023], i.e., satisfies the error guarantee even if built by an arbitrary sequence of merge operations.
Theorem 3.
Suppose that we build a SplineSketch using any sequence of pairwise merge operations executed on $n$ data items. Then SplineSketch has worst-case rank error of $O(\log(\alpha)\cdot n/k)$.
Proof.
First, we assume w.l.o.g. that there are no items of frequency above $n/k$ by running a Misra-Gries sketch [Misra and Gries, 1982] of size $O(k)$ in parallel, as described in Sec. 3.3. We note that the Misra-Gries sketch is fully mergeable [Agarwal et al., 2013] and that it has error at most $n/k$ for estimating the frequency of any item stored in the sketch. Second, we show an upper bound on the error only for the thresholds of the final sketch. The bound then follows for all other items by the fact that each bucket has $O(n/k)$ items. Thus, in the following we consider a threshold $\tau_{i}$ in the final sketch created by merge operations from $n$ items in an arbitrary way, and analyze the rank error at $\tau_{i}$. We consider the binary tree $T_{0}$ corresponding to merging sketches and operations with buckets described in Section B.1.
The error at threshold $\tau_{i}$ is increased either due to splitting a bucket or due to taking a union of buckets when merging sketches. We first analyze the error from the splits. For the threshold $\tau_{i}$, we consider the following sequence of tree nodes $t_{0},t_{1},t_{2},\dots$ and thresholds $\sigma_{0},\sigma_{1},\sigma_{2},\dots$: Let $\sigma_{0}:=\tau_{i}$ and let $t_{0}$ be the split-node such that the split creates $\tau_{i}$ and there is no node on the path from $t_{0}$ to the root that removes $\tau_{i}$ by a join of buckets. Note that the rank error at $\tau_{i}$ will not increase after the split represented by $t_{0}$, by the description of the algorithm. Then for $a=1,2,\dots$, we define $\sigma_{a}$ and $t_{a}$ as follows: If $\sigma_{a-1}$ is created by a split of a bucket, consider the split of a bucket $(\tau^{\prime},\tau^{\prime\prime}]$ that creates $\sigma_{a-1}=(\tau^{\prime}+\tau^{\prime\prime})/2$ during the consolidate method at $t_{a-1}$. We set $\sigma_{a}:=\tau^{\prime}$ and let $t_{a}$ be the split-node in the subtree of $t_{a-1}$ such that the split at $t_{a}$ creates $\sigma_{a}$ and $t_{a}$ is the closest node to $t_{a-1}$ with the property that on the path from $t_{a}$ to $t_{a-1}$, there is no join of buckets that removes $\sigma_{a}$. Otherwise, $\sigma_{a-1}$ is created by evenly spaced selection of initial buckets at $t_{a-1}$ (i.e., the sketches represented in the subtree of $t_{a-1}$ have no buckets, except for $t_{a-1}$), and we stop this process. Let $m$ be the index when this process stops; that is, at $t_{m}$, threshold $\sigma_{m}$ was created by the evenly spaced selection.
To analyze the rank error, for a threshold $\tau$ in the sketch represented by a tree node $t$, we define $\operatorname{Err}^{t}(\tau)$ as the rank error of $\tau$ in the sketch represented by $t$, i.e., the absolute value of the difference between the rank of $\tau$ among items at leaves in the subtree of $t$ (i.e., the items summarized by the sketch at $t$) and the estimated rank of $\tau$, which is the prefix sum of counters of buckets up to $\tau$. We have that $\operatorname{Err}^{t_{m}}(\sigma_{m})=0$, as the evenly space selection is exact at the bucket thresholds, and that $\operatorname{Err}^{t_{0}}(\sigma_{0})=\operatorname{Err}^{t_{0}}(\tau_{i})=\operatorname{Err}^{t_{0}}(r)$, where $r$ is the root, representing the final sketch.
Let $j_{a}$ be the epoch of the sketch represented by $t_{a}$, i.e., $t_{a}$ is a $j_{a}$-node. We claim that for any epoch $j$, there are at most $O(\log(\alpha))$ nodes $t_{a}$ with $j_{a}=j$. Indeed, consider the $j$-nodes on the path $P_{j}$ from the lowest $j$-node to the root of the tree. Since any threshold created by a split in a $j$-node is not removed by a join in a $j$-node due to the protection, any threshold $\sigma_{a}$ created by a $j$-node $t_{a}$ can only be removed by a join of buckets in a $j^{\prime}$-node for $j^{\prime}>j$. Let $b$ be the minimum $a$ such that $j_{a}=j$, i.e., $\sigma_{b}$ is the last threshold created by a $j$-node. Note that after the split in $j$-node $t_{a}$ creating $\sigma_{a}$, the bucket $(\sigma_{a},\sigma^{\prime}_{a}]$ contains $\sigma_{b}$ and as the next threshold $\sigma^{\prime}_{a}$ gets protected, for any $d\in(b,a]$, it holds that $\sigma_{b}\in(\sigma_{d},\sigma^{\prime}_{d}]$ Observe that by the definition of $\alpha$ and since we always split a bucket in the middle, after $O(\log(\alpha))$ splits in $j$-nodes on path $P_{j}$ that involve buckets containing $\sigma_{b}$, the length of the bucket containing $\sigma_{b}$ will be below the smallest distance between distinct items, i.e., the bucket will cover at most one distinct item $x$ of the input. The algorithm may still split the bucket containing $\sigma_{b}$ after the $O(\log(\alpha))$ splits but only $O(1)$ times, because every split decreases the bucket counter by a factor of less than $1-\beta\approx 0.934$, we do not split a bucket with counter below $(C_{b}/100)\cdot n/k$ (even due to the heuristic error), and there are no items of frequency above $n/k$ by the assumption. This proves the claim that there are at most $O(\log(\alpha))$ nodes $t_{a}$ with $j_{a}=j$.
Moreover, each node $t_{a}$ with $j_{a}=j$ increases the error by at most $O(n_{j}/k)$, due to bucket size bounds; specifically for any $a$ with $j_{a}=j$, we have that $\operatorname{Err}^{t_{a}}(\sigma_{a})\leq\operatorname{Err}^{t_{a+1}}(\sigma_{a+1})+O(n_{j}/k)$. Therefore, the total error from splits in $j$-nodes is at most $O(\log\alpha\cdot n_{j}/k)$. Since the values of $n_{j}$ increase geometrically, the total error at $\tau_{i}$ resulting from splits sums to $O(\log\alpha\cdot n/k)$.
To analyze the error from taking the union of buckets in merge-nodes of $T_{0}$, consider a merge-node $t$ such that $\sigma_{a}$ is a threshold of the resulting bucket. Observe that the additional error incurred due to taking the union of buckets is $1/k$ times the number of items summarized by source sketch $S^{\prime}_{t}$ of the merge operation that does not contain $\sigma_{a}$ (if both source sketches contain $\sigma_{a}$, we choose arbitrarily which one is discarded). Further, observe that for any epoch $j$, the total size of such sketches affecting the error at $\sigma_{a}$ with $j_{a}=j$ is at most $O(n_{j})$, since otherwise, the epoch would switch to $j^{\prime}>j$. Since the $n_{j}$’s form a geometric sequence, the total error at $\tau_{i}$ from all merge-nodes is at most $O(n/k)$. Hence, $\operatorname{Err}^{t_{0}}(r)\leq O(\log\alpha\cdot n/k)$. ∎
Appendix C Reproducibility of Experimental Evaluation
Here, we provide more detailed information on running our benchmarks. The code to run the experiments, including prototypes of SplineSketch in Python and Java and code to generate the synthetic datasets, is available at https://github.com/PavelVesely/SplineSketch-experiments. Experiments were performed on an AMD EPYC 7302 (3 GHz) server with 251 GB RAM and SSD storage.
Quantile sketches.
We have used the Java prototype of SplineSketch. We have also evaluated $t$-digest [Dunning and Ertl, 2019] (v3.3, https://github.com/tdunning/t-digest), MomentSketch [Gan et al., 2018] (https://github.com/stanford-futuredata/msketch), and KLL [Karnin et al., 2016] (Java implementation by Apache DataSketches, v6.0.0, https://github.com/apache/datasketches-java). For $t$-digest, we use the default merging variant with $k_{0}$ scale function that aims at the uniform error [Dunning and Ertl, 2019]. KLL was used in the variant with the double data type, i.e., an instance of KllDoublesSketch. We note that the MomentSketch implementation only answers quantile queries, while our experiments required rank estimates; we have simulated rank queries using binary search that was done for all $m$ queries in parallel in order to decrease the number of calls to the quantile query method from $\approx m\log_{2}n$ (querying just one rank) to $\approx\log_{2}n$ calls (querying $m$ ranks), to avoid repeated computation of the distribution fitting the moments and log-moments (which is a costly operation).
The sketch size is measured in bytes when serialized on disk for storage without supporting data structures such as buffers. More precisely, for SplineSketch and $t$-digest we count 16 bytes per bucket/centroid, for MomentSketch we measured $1616k+$ 16 bytes ($k$ moments and log-moments, minimum, and maximum), and for KLL we count 8 bytes per stored item.
Datasets.
[…]
Rank queries are generated by equally spaced selection from sorted data. The queries are executed in one batch whenever possible.
Appendix D Figures with Complete Experimental Evaluation
We provide additional figures with experimental results. Figs. 6 and 7 contain the update and query times depending on the sketch size. Fig. 8 and 9 contains complete results for the average and maximum error, respectively.
Figure 6. Time per update in $\mu$s depending on sketch size on $n=10^{8}$ samples from normal distribution.
Figure 7. Time per query in $\mu$s depending on sketch size for $10^{5}$ queries on $n=10^{8}$ samples from normal distribution.
(a) HEPMASS dataset [Whiteson, 2016]
(b) Power dataset [Hebrail and Berard, 2006]
(c) Books dataset [Kipf et al., 2019, Marcus et al., 2020]
(d) Normal distribution
(e) Uniform distribution
(f) Pareto distribution
(g) Lognormal distribution
(h) Loguniform distribution
(i) Signed loguniform distribution with large exponents
(j) Normal distribution with a small parameter change after $n/2$ items
(k) Normal distribution with a large parameter change after $n/2$ items
(l) Normal distribution with $n/2$ items, followed by $n/2$ high-frequency items
Figure 8. Average rank error (log-scale) depending on the sketch size in bytes. The average is taken over $10^{5}$ evenly spaced queries.
(a) HEPMASS dataset [Whiteson, 2016]
(b) Power dataset [Hebrail and Berard, 2006]
(c) Books dataset [Kipf et al., 2019, Marcus et al., 2020]
(d) Normal distribution
(e) Uniform distribution
(f) Pareto distribution
(g) Lognormal distribution
(h) Loguniform distribution
(i) Signed loguniform distribution with large exponentslabel
(j) Normal distribution with a small parameter change after $n/2$ items
(k) Normal distribution with a large parameter change after $n/2$ items
(l) Normal distribution with $n/2$ items, followed by $n/2$ high-frequency items
Figure 9. Maximum rank error (log-scale) depending on the sketch size in bytes. The maximum is taken over $10^{5}$ evenly spaced queries.