### v2

# SplineSketch: Even More Accurate Quantiles with Error Guarantees

Abstract.
Space-efficient streaming estimation of quantiles in massive datasets is a fundamental problem with numerous applications in data monitoring and analysis. While theoretical research led to optimal algorithms, such as the Greenwald-Khanna algorithm or the KLL sketch, practitioners often use other sketches that perform significantly better in practice but lack theoretical guarantees. Most notably, the widely used $t$-digest has unbounded worst-case error.
In this paper, we seek to get the best of both worlds. We present a new quantile summary, SplineSketch, for numeric data, offering near-optimal theoretical guarantees, namely uniformly bounded rank error, and outperforming $t$-digest by a factor of 2-20 on a range of synthetic and real-world datasets. To achieve such performance, we develop a novel approach that maintains a dynamic subdivision of the input range into buckets while fitting the input distribution using monotone cubic spline interpolation. The core challenge is implementing this method in a space-efficient manner while ensuring strong worst-case guarantees.
quantile estimation, streaming algorithms, mergeable summaries, data sketches, cubic spline interpolation
journalyear: 2025conference: ; ; copyright: noneccs: Theory of computation Sketching and samplingccs: Theory of computation Streaming modelsccs: Theory of computation Data structures and algorithms for data management
1. Introduction
Data sketching has become one of the main tools for dealing with massive data volumes. Sketches provide a scalable way to extract key features from large datasets, enabling real-time analysis at streaming speed or massively parallel processing of distributed datasets, in applications such as network monitoring, machine learning, privacy, or bioinformatics.
Approximating order statistics is a central sketching problem. The goal is to process a massive dataset that is possibly distributed and summarize it in a small space into a sketch that is updatable with new observations, mergeable, and provides highly accurate estimates for the median, percentiles, and their generalization, quantiles. Equivalently, we can estimate the data distribution by approximating ranks, i.e., the number of input items no larger than a given element. Quantile sketches have been applied, e.g., to monitoring latencies (Tene, 2015) or for detecting anomalies (Siffer et al., 2017).
Streaming quantile estimation has thus received a significant attention in research community. From the theory point of view, it has been solved to a large extent. Namely, optimal algorithms are known for the uniform error, such as the Greenwald-Khanna (GK) sketch (Greenwald and Khanna, 2001; Cormode and Veselý, 2020) and KLL (Karnin et al., 2016). However, the accuracy of these theory-based quantile summaries on real-world data is only a small constant factor better than the worst-case bound. Instead, practitioners often use algorithms that perform significantly better in practice but have no guarantees. Most notably, $t$-digest (Dunning and Ertl, 2019) is widely used and, according to (Dunning, 2021), it has been adopted by major tech companies (reportedly Microsoft, Facebook, Google) and open-source projects (Elasticsearch, TimescaleDB, Apache Dubbo, etc.). At the same time, no matter how large the $t$-digest is, its error can be arbitrarily bad (Cormode et al., 2021). Furthermore, as we outline in Sec. 2, to the best of our knowledge, other quantile sketches do not provide both worst-case bounds and error similar to $t$-digest in practice.
Here, we introduce a new quantile sketch, called SplineSketch, that not only gets the best of both worlds – theoretical guarantees and very high accuracy in practice – but also significantly outperforms the state of the art, including $t$-digest (Dunning and Ertl, 2019). Our sketch processes any numerical input in a streaming fashion and is fully mergeable, enabling efficient parallel or distributed processing of large datasets. Before explaining our contribution in more detail, we define our setting, including the considered error guarantees.
Definitions.
We use $n$ for the number of items currently summarized by the sketch. For a multiset $S$ of real numbers and any $x\in\mathbb{R}$, let $\mathrm{rank}(x)=|\{y\in S:y\leq x\}|$, i.e., the number of input items smaller than or equal to $x$. For any $q\in(0,1]$, we call $y:=\min\{x\in S:\mathrm{rank}(x)\geq q\cdot n\}$ the $q$-quantile of $S$, i.e., the $q$-quantile is the $\lceil q\cdot n\rceil$-th largest item in $S$ (if there are no duplicates). This definition implies that a method for estimating ranks may be used for estimating quantiles with the same accuracy, by performing binary search. The aspect ratio $\alpha$ of $S$ is defined as $\alpha=(\max(S)-\min(S))/\gamma$, where $\gamma=\min_{i\neq j:x_{i}\neq x_{j}}|x_{i}-x_{j}|$ is the smallest difference of distinct items on input, assuming at least two distinct items are present.
We focus on the streaming and mergeability settings. The streaming model requires that the dataset $S$ is processed in one pass with low memory (in an arbitrary order, possibly adversarial). In the more general setting of mergeable summaries (Agarwal et al., 2013), $S$ is arbitrarily split into a number of parts, each summarized in the streaming fashion and the resulting summaries are combined via pairwise merge operations in an arbitrary way. In other words, besides answering rank and quantile queries, the sketch needs to have an update operation, which adds a new observation to the dataset, and a merge operation, which takes two sketches and outputs a single sketch of their combined input.
Following a long line of work, e.g., (Shrivastava et al., 2004; Greenwald and Khanna, 2001; Agarwal et al., 2013; Karnin et al., 2016; Luo et al., 2016; Cormode and Veselý, 2020; Assadi et al., 2023; Gupta et al., 2024; Gribelyuk et al., 2024), we study the uniform rank error which for a given accuracy parameter $\varepsilon>0$ and any query item $x$ (that may or may not be in the input) requires the algorithm to return a rank estimate $\widehat{\mathrm{rank}}(y)$ such that $|\widehat{\mathrm{rank}}(x)-\mathrm{rank}(x)|\leq\varepsilon\cdot n$, and for any quantile query $q\in(0,1]$, the algorithm should return a $q^{\prime}$-quantile for $q^{\prime}=q\pm\varepsilon$.
(a) HEPMASS dataset (Whiteson, 2016)
(b) Power dataset (Hebrail and Berard, 2006)
(c) Books dataset (Kipf et al., 2019; Marcus et al., 2020)
Figure 1. Average error of SplineSketch, KLL, MomentSketch, $t$-digest, and GKAdaptive on three real-world datasets (Sec. 5).
Our contributions
Our main contribution is a new technique for designing quantile summaries with bounded worst-case error and high accuracy in practice, combining dynamically maintained subdivision of the input range with a suitable interpolation over the buckets, outlined in Section 1.1 and described in Section 3.
We demonstrate the power of this technique on a new quantile sketch, called SplineSketch. It requires a single parameter $k$ that determines its space usage, which is $O(k)$ memory words for processing the input and $16\cdot k$ bytes when serialized for storage. The sketch deterministically guarantees a theoretically near-optimal uniform rank error on any input, close to the best possible error of $O(n/k)$ with $k$ memory words; namely, the error is $O(n/k\cdot\log\alpha)$, where $\alpha$ is the aspect ratio. At the same time, the accuracy of our sketch in practice is in fact typically 10–200 times better than $n/k$. Furthermore, we prove our sketch is fully mergeable while retaining the guarantees, i.e., we analyze it in the most general setting when the sketch is constructed by an arbitrary sequence of pairwise merge operations. Finally, our sketch can be resized based on the current memory requirements.
[…]
Overall, we demonstrate that besides mergeability, worst-case guarantees, and high accuracy in practice, our approach offers flexibility, making it suitable for a range applications.
1.1. Overview of our approach
The core of our approach is in maintaining a subdivision of the input range into buckets in a way that better adapts to the input distribution. Namely, SplineSketch consists of thresholds $\tau_{1},\cdots,\tau_{k}$ and then $(\tau_{i-1},\tau_{i}]$ is the $i$-th bucket, with $\tau_{0}=$-$\infty$. For each bucket, we store a counter $b_{i}$ that represents an approximate number of items that lie in $(\tau_{i-1},\tau_{i}]$; we refer to $b_{i}$ as the size of the bucket $i$. Furthermore, we use cubic spline interpolation in order to answer queries more accurately in practice; see Fig. 2 for an illustration.
Figure 2. SplineSketch illustration, with the values and thresholds $\tau_{i}$ on the $x$ axis and the rank space $[0,n]$ on the $y$ axis. The true cumulative distribution function of the data is depicted as a red dashed curve. The green dots correspond to the prefix sums of bucket counters, i.e., the counter $b_{i}$ is the difference of the $y$ values of the green dots at $\tau_{i}$ and $\tau_{i-1}$. The solid black curve is an interpolation over the green dots.
Answering queries using cubic splines.
The estimated rank of any threshold $\tau_{i}$ is $\sum_{j=1}^{i}b_{i}$. It is however not clear how to answer rank queries that are not at one of the bucket boundaries $\tau_{i}$. Most of the previous work would just instead query the closest value $\tau_{j}$ and return that as the answer, or perform a linear interpolation. However, we seek to do better than that, using more advanced interpolation methods, namely the piecewise cubic Hermite interpolating polynomial (PCHIP) interpolation (Fritsch and Butland, 1984; Fritsch and Carlson, 1980). Specifically, if we are querying $x$, such that $\tau_{i}<x<\tau_{i+1}$, we estimate the ranks $\hat{r}_{i-1},\cdots,\hat{r}_{i+2}$ of $\tau_{i-1},\cdots,\tau_{i+2}$, then compute the PCHIP interpolation in order to interpolate between $\hat{r}_{i}$ and $\hat{r}_{i+1}$, and evaluate this polynomial at $x$, which is our estimate.
Maintaining the buckets.
The most important choice we need to make is what the thresholds of the buckets should be. We always set $\tau_{1}$ and $\tau_{k}$ to be the smallest and the largest item, respectively, that we have seen so far. Other than that, as the new data keeps arriving, we will be adjusting the buckets by splitting some of them and joining some pairs of adjacent ones. When we join two adjacent buckets, we simply remove the threshold that they share and set the new counter to the sum of the two buckets’ counters. In the split procedure, we split the bucket in the middle. We then need to choose how to split the count into the two new sub-buckets. To this end, we perform a query at the midpoint (as described above) and we divide the count so as to make the value at the midpoint as returned by the query. It remains to specify how we choose which buckets to split and join. In the description below, we ignore the efficiency. In the end, we use a buffer and perform updates in batches, obtaining low amortized update time.
There are two rules for choosing a bucket to split. Firstly, throughout the execution, we ensure that all buckets have counter $O(n/k)$, and we split a bucket when it gets too big (the exact value of the bucket bound is defined in Sec. 3.1). We later use these size bounds, together with other invariants, to show the worst-case error bounds. Intuitively speaking, we show that no item can be part of too many splits and that each split can only contribute to the error by at most the size of the bucket the item belongs to.
The second rule for bucket splitting is intended to get significantly better error in practice, while preserving the good error bounds in the worst case. To this end, we also introduce for each bucket a number which we call heuristic error, which is our guess for the actual error of the cubic spline inside the bucket. The underlying reason is that how the data distribution function changes is captured by the derivatives of the distribution. To illustrate the intuition behind this notion, suppose the data items are i.i.d. from an unknown distribution. If the growth of its CDF $F$ does not change too much around a bucket, specifically the absolute value of its second derivative $F^{\prime\prime}$ is small, the error of the spline approximation should be low as well. In practice, we do not know the CDF and the data may not be i.i.d., but we nevertheless use an estimate of the CDF’s second derivative around a bucket as a heuristic approximation of the true error. We do not store this number and instead compute it when needed. If the maximum heuristic error can be decreased substantially by splitting one bucket and joining two other adjacent buckets, while preserving bucket bounds, then we perform the split and join of these buckets. The exact function that we use as the heuristic error is in Eqn. 2 in Sec. 3.1.
To ensure we always have exactly $k$ buckets, whenever we split a bucket, we join some pair of adjacent buckets. When we perform a join, we would like to choose two adjacent buckets which together would form a bucket with the lowest possible value of the heuristic error. However, we need to ensure that the new bucket would not grow too large. Therefore, we only consider pairs for which the size of the resulting bucket would be not too close to the bucket bound, specifically at most 75% of the bucket bound.
Finally, we do not want to perform too many joins and splits on a single bucket, as it causes the error to grow. To this end, we divide the input into epochs, where one epoch is defined as time during which the input size increased by a fixed constant factor. If the bucket was created by a split, we mark its thresholds as protected and we do not allow it to be joined during the same epoch. We prove that despite all these constraints after a split there will be always at least one pair of adjacent buckets that can be joined.
Merging two sketches.
The merge operation for our sketch is conceptually simple: Given two SplineSketches on input, we first create a combined set of thresholds by taking the union of both sketches’ thresholds and removing duplicates. We then assign counters to the new buckets based on the sum of their estimated ranks in both original sketches, for which we use precomputed interpolations over the input sketches. Taking the union of thresholds typically results in a sketch with more buckets than desired, so we reduce the number of buckets by joining adjacent ones with the lowest heuristic error, until we reach the target bucket count. To maintain worst-case guarantees, if the two sketches being merged differ significantly in size, the threshold protection from the larger sketch is preserved during merging; however, when their size is similar, we basically reset the epoch, together with the threshold protection.
Resizing the sketch.
Using joining or splitting the buckets allows us to easily change the number $k$ of buckets. Namely, when we want to make the sketch smaller, we just join an appropriate number of adjacent buckets without any splits, whereas for increasing the sketch size, we perform splits without joins. This demonstrates the flexibility of our approach, and can be useful in practice, e.g., when we allocate more memory for processing the input and then compress the sketch into much smaller size for storage, which leads to overall better accuracy than using the smaller size for processing.
Paper outline.
Section 2 outlines the most relevant work on quantile sketches and provides a comparison with our approach. The full algorithmic description of SplineSketch appears in Section 3, together with its theoretical properties. The implementation details appear in Section 4 and the experimental evaluation in Section 5. We conclude with discussing the pros and cons of our approach and the open problems in Section 6.
2. Related work
Quantile summaries are intensively investigated since the 90s, with the pioneering work of Munro and Paterson from 1978 (Munro and Paterson, 1980) who showed that exact quantile selection in the streaming setting, i.e., one pass with sublinear space, is impossible. In the comparison-based model, Greenwald and Khanna (Greenwald and Khanna, 2001) designed a deterministic algorithm that stores $O(\varepsilon^{-1}\cdot\log\varepsilon n)$ stream items and guarantees $\pm\varepsilon\cdot n$ uniform error; this bound is optimal among deterministic comparison-based quantile summaries (Cormode and Veselý, 2020). However, the GK summary is not known to be mergeable while retaining its space bounds (Agarwal et al., 2013), despite that the intricate analysis from (Greenwald and Khanna, 2001) was recently simplified and generalized to weighted updates (Assadi et al., 2023; Gribelyuk et al., 2024). Randomization allows to remove the logarithmic dependency on the stream length $n$, and after a sequence of improvements (Manku et al., 1999; Luo et al., 2016; Felber and Ostrovsky, 2015), Karnin, Lang, and Liberty (Karnin et al., 2016) developed an optimal randomized algorithm that stores just $O(\varepsilon^{-1})$ items (with a constant probability of a too large error). Beyond the comparison-based model, $q$-digest (Shrivastava et al., 2004) uses an implicit binary tree over the items’ universe $\mathcal{U}$ and stores a subset of the tree nodes with associated counters. It requires space of $O(\varepsilon^{-1}\cdot\log|\mathcal{U}|)$ and a foreknowledge of $\mathcal{U}$. Very recently, Gupta, Singhal, and Wu developed a compressed variant of $q$-digest that fits into $O(1/\varepsilon)$ memory words, by packing more nodes into one memory word (Gupta et al., 2024). However, unlike SplineSketch, all of these algorithms do not offer practical accuracy on real-world data that goes beyond their worst-case behavior, as demonstrated in Section 5.
Another line of work focused on the stronger relative (multiplicative) error guarantee, which requires that the error for an item of rank $r$ is at most $\pm\varepsilon\cdot r$. For deterministic comparison-based algorithms, there is a gap of $O(\log\varepsilon n)$ between the merge-and-prune algorithm using space $O(\varepsilon^{-1}\cdot\log^{3}\varepsilon n)$ (Zhang and Wang, 2007) and the lower bound of $\Omega(\varepsilon^{-1}\cdot\log^{2}\varepsilon n)$ (Cormode and Veselý, 2020). Efficient randomized sketches with relative error have only appeared recently, namely, ReqSketch with space $O(\varepsilon^{-1}\cdot\log^{1.5}\varepsilon n)$ (Cormode et al., 2023) and its “elastic“ version achieving space close the information-theoretic lower bound of $\Omega(\varepsilon^{-1}\cdot\log\varepsilon n)$ (Gribelyuk et al., 2025); however, mergeability of the “elastic” version is open, while ReqSketch is analyzed also in the mergeability setting. There is also an extension of $q$-digest to the relative error that uses space $O(\varepsilon^{-1}\cdot\log\varepsilon n\cdot\log|\mathcal{U}|)$ (Cormode et al., 2006); it is not known to be optimal. Compared to SplineSketch, the main benefit of relative error is better accuracy for the tails which, however, comes at the cost of higher space complexity, both in theory (additional dependency on $\log n$) and practice (see e.g. (Cormode et al., 2021)). Thus, relative-error sketches are incomparable to SplineSketch and other uniform-error algorithms. Finally, we note that even stronger error guarantees, such as the q-error as defined in (Moerkotte et al., 2014), cannot be achieved in the streaming or mergeability settings with sketches of sublinear size, due to standard space lower bounds from the one-way communication complexity of indexing.
For dynamic streams (i.e., with deletions), one can only achieve the uniform error guarantee in sublinear space, and the best quantile summary supporting deletions is Dyadic CountSketch (Luo et al., 2016). The support for deletions comes at an additional cost of a polylogarithmic dependency on the universe size $|\mathcal{U}|$, while insertion-only sketches, including KLL or SplineSketch, are much more space-efficient.
In contrast to rank error, DDSketch (Masson et al., 2019) and UDDSketch (Epicoco et al., 2020) provide the relative value error (i.e., an item close to the desired quantile in the item space) based on maintaining a suitable exponential histogram. While this usually captures the distribution tails well, the rank error can be arbitrarily large as these sketches in fact fail to capture most of the distribution in sufficient detail, leading to overall worse accuracy than KLL, $t$-digest, or SplineSketch. Furthermore, the value error is not invariant to natural data transformations such as translations.
Finally, there are more heuristic approaches that, similarly as SplineSketch, work only for numerical data, but do not aim for any worst-case rank guarantees, unlike SplineSketch. Specifically, $t$-digest (Dunning and Ertl, 2019; Dunning, 2021) performs an online one-dimensional $k$-means clustering by averaging data items into a given number of centroids, and interpolates linearly between the centroids. A downside of this approach is that many of the items summarized by a centroid may be smaller or larger than an adjacent centroid mean, implying under- or over-estimation. That is, the centroids are only ordered by their means, while the represented items may be far from ordered. In fact, one can impose high overlaps of represented items, leading to almost arbitrarily large error on adversarial instances (Cormode et al., 2021). Linear interpolation techniques were combined with the KLL sketch to improve its accuracy (Schiefer et al., 2023).
Finally, MomentSketch (Gan et al., 2018; Mitchell et al., 2021) is possibly the most compact summary as it consists of $k$ moments and log-moments of data items, for a small $k$ such as $k=15$. Upon a query, the sketch constructs a distribution with the same moments and log-moments, using the maximum entropy principle and Chebyshev polynomials, which is a costly operation. However, MomentSketch only works well for certain smooth distributions but, as our experiments also demonstrate, suffers a large error in many real-world cases, including the uniform distribution.
The aforementioned quantile sketches form a basis for per-key quantile estimation (a simultaneous estimation of quantiles for multiple substreams in the main stream), as investigated in a recent line of work (Dong et al., 2024; Guo et al., 2023; He et al., 2023; Shahout et al., 2023).
Additionally, we remark that the idea of approximating the data distribution using splines is not new and has been widely used in computer science before, e.g., for query optimization (Neumann and Michel, 2008) and indexing (Kipf et al., 2020) in database systems.
3. Description of SplineSketch
We provide a description of our algorithm, omitting implementation details, explained in Sec. 4. In fact, there are many possible implementations of the sketch that still satisfy the worst-case guarantee. Our sketch works for any numerical input, i.e., items are integers or floating-point numbers. We first present the sketch assuming there are no items of high frequency, namely appearing more than (n1)/(k1)(n-1)/(k-1) times in the input after inserting any number $n$ of items into SplineSketch of size $k$. We show how to remove this assumption in Sec. 3.3.
Buckets.
Fix an integer parameter $k\geq 6$ (as required in the analysis). The sketch primarily consists of $k$ distinct thresholds $\tau_{1}<\tau_{2}<\cdots<\tau_{k}$ and $k$ counters $b_{1},\ldots,b_{k}$, where $b_{i}$ is our estimate of the number of items that lie in the interval $(\tau_{i-1},\tau_{i}]$; we define $\tau_{0}=$-$\infty$. We call these intervals buckets and refer to a bucket by the index of its right threshold, so $(\tau_{i-1},\tau_{i}]$ is called ”bucket ii” or the $i$-th bucket. We also frequently refer to $b_{i}$ as the size of bucket $i$ and we define the length of bucket $i$ as $\ell_{i}:=\tau_{i}-\tau_{i-1}$.
Throughout the execution of the algorithm $\tau_{1}$ and $\tau_{k}$ are always set to be the minimum and the maximum item in the stream, respectively. This makes the first bucket special as its counter only stores the frequency of $\tau_{1}$.
We also maintain a buffer of size $O(k)$ that we use for faster processing of incoming items, i.e., low average update time, and one auxiliary bit-array of length $k$ for “protected thresholds” (Sec. 3.1). After processing the whole input, the buffer is merged into the buckets and the auxiliary bit-array can be discarded. We maintain two invariants: All bucket counters plus the number of items in the buffer sum up to $n$ and no bucket is empty.
Insertion and initial buckets.
When a new element arrives, we first put it in the buffer. When the buffer gets full, we execute the consolidate method (Sec. 3.1) that may change bucket thresholds and merges the buffer into buckets. The only exception is the first time when the buffer gets full. In that case, we simply sort it, pick the first and then every $\approx(n-1)/(k-1)$-th item, selecting $k$ items as the initial thresholds, and set counters $b_{i}$ to exactly the number of items in the corresponding buckets. Since no item has frequency greater than (n1)/(k1)(n-1)/(k-1), these selected items are distinct and every bucket is nonempty. For the pseudocode of the bucket initialization, see Algorithm 1.
Query.
For a rank query at a threshold $\tau_{i}$, we return the sum of the counters up to the $i$-th, i.e., $\sum_{j=1}^{i}b_{i}$. For $x$ strictly inside the $i$-th bucket, i.e. $\tau_{i-1}<x<\tau_{i}$, we use an interpolation method to estimate the rank of $x$; note that this works no matter whether $x$ appeared on input or not. The purpose of using a suitable interpolation is to get much better error in practice, while not affecting the worst-case bounds (see Sec. 3.3).
Specifically, we use a monotone variant of the Piecewise Cubic Hermite Interpolating Polynomial (PCHIP) (Fritsch and Butland, 1984; Fritsch and Carlson, 1980), which is an interpolation method that constructs a continuous function by fitting cubic polynomials between data points. Unlike standard cubic spline interpolation, PCHIP adjusts the first derivatives at each data point to preserve the shape and monotonicity of the original data. For an introduction in greater detail, including the formulas for the interpolation, see (Wong, 2020). Since the ranks of the input items are non-decreasing, preservation of monotonicity makes PCHIP suitable for our application.
We remark that computing this interpolation requires only the knowledge of $\tau_{i-2},\cdots,\tau_{i+1}$ together with corresponding prefix sums of bucket counters. Therefore, provided that we maintain the prefix sums, the computation of the interpolation takes $O(1)$ time. The interpolation does not extrapolate, i.e., for $x<\tau_{1}$, it returns 0, and for $x>\tau_{k}$, it returns $\sum_{j=1}^{k}b_{i}$. Since one has to find the right bucket by binary search, one query takes time $O(\log k)$ provided that the prefix sums are precomputed.
In principle, our approach can be used with any interpolation method (ideally one that preserves monotonicity). One popular alternative would be to use kernel smoothers. However, this particular method would introduce additional error on the thresholds, which would be harder to control. We leave the exploration of alternative interpolation methods for future work. For an extensive treatment of different interpolation methods, see (Fan and Gijbels, 1996).
3.1. Consolidating buckets
The main procedure of the sketch is Consolidate. The method first merges the buffer into the current buckets, by counting the new number of items in each bucket, and updates the counters accordingly. Then, it keeps changing the buckets based on the new values of the counters, possibly removing some of the thresholds and creating new ones. Finally, the buffer gets emptied. The crux of the consolidate method is how to modify the thresholds, which we describe in the remainder of this subsection. For the pseudocode of Consolidate, see Algorithm 2.
Figure 3. Illustration of joining a pair of buckets and splitting one bucket (depicted similarly as in Fig. 2). The join removes threshold $\tau_{j}$ and merges (sums up) the two counters of the joined buckets. The split takes the $i$-th bucket $(\tau_{i},\tau_{i+1}]$ and inserts a new threshold $\tau^{\prime}_{i+1}=(\tau_{i}+\tau_{i+1})/2$ in the middle, setting the counters of the two resulting buckets $(\tau_{i},\tau^{\prime}_{i+1}]$ and $(\tau^{\prime}_{i+1},\tau_{i+1}]$ using the interpolation (solid black curve). The thresholds $\tau_{i},\tau^{\prime}_{i+1}$, and $\tau_{i+1}$ get protected from joining, as depicted by the locks, and this protection is removed at the end of the epoch.
Joining and splitting buckets.
We have two basic operations that change bucket thresholds, join and split (Fig. 3). A join takes two neighboring buckets $(\tau_{i-1},\tau_{i}]$ and $(\tau_{i},\tau_{i+1}]$ and turns them into a single bucket $(\tau_{i-1},\tau_{i+1}]$ with its counter set to $b_{i}+b_{i+1}$. Note that after joining, buckets are (implicitly) renumbered. That is, we actually set new thresholds $\tau^{\prime}$ and counters $b^{\prime}$ as $\tau^{\prime}_{i}=\tau_{i+1}$, $b^{\prime}_{i}=b_{i}+b_{i+1}$, and for any $j>i$, $\tau^{\prime}_{j}=\tau_{j+1}$ and $b^{\prime}_{j}=b_{j+1}$. The buckets up to $\tau_{i-1}$ do not change.
When splitting bucket $(\tau_{i-1},\tau_{i}]$ for $i>1$, we replace it by two buckets $(\tau_{i-1},\frac{\tau_{i-1}+\tau_{i}}{2}]$ and $(\frac{\tau_{i-1}+\tau_{i}}{2},\tau_{i}]$. The bucket counters are set as follows: We use the query method to get estimated ranks $\hat{r}$ at $\tau_{i-1},\frac{\tau_{i-1}+\tau_{i}}{2},\tau_{i}$. We then set the counters for the two new buckets to $\hat{r}(\frac{\tau_{i-1}+\tau_{i}}{2})-\hat{r}(\tau_{i-1})$ and $\hat{r}(\tau_{i})-\hat{r}(\frac{\tau_{i-1}+\tau_{i}}{2})$, respectively. Other buckets are again implicitly renumbered.
Observe that every split increases and every join decreases the number of buckets by one. Thus, after every split we join some other pair of buckets, which ensures there are always $k$ buckets (unless a change of $k$ is desired). The crux of the algorithm is in selecting which buckets to split and join, which we describe below.
Protected thresholds and epochs.
After splitting a bucket, the estimated rank, and therefore the estimation error, at its right threshold stays the same. However, the error at the newly created threshold may be bigger, in the worst case by the size of the original bucket counter. If we repeatedly join and split the newly created buckets with their neighbors, we might let the error increase too much. With the small overhead of $k$ bits when processing the input, we show how to avoid too frequent splits and joins at a single location and prevent the associated error accumulation. To this end we introduce the notion of protected thresholds.
Specifically, we divide the input into epochs. The first epoch has length $\Theta(k)$ and each successive epoch starting at time $t$ has length $0.25\cdot t$. When we split a bucket $(\tau_{i-1},\tau_{i}]$, we mark all three thresholds $\tau_{i-1},\tau^{\prime}_{i},\tau_{i}$ as protected, where $\tau^{\prime}_{i}=(\tau_{i-1}+\tau_{i})/2$ is the new threshold. We store a bit vector of length $k$, with the $i$-th bit set to 1 when the $i$-th threshold is protected. When a new epoch starts, we reset all thresholds to the default unprotected state by zeroing the bit vector. As specified later in Def. 1, we never join buckets that have a protected threshold between them.
Note that we may still split a single bucket multiple times during one epoch, or even during one call of Consolidate, namely, if the bucket receives many new items. Similarly, a single bucket may be joined multiple times during one consolidation, but a combination of splitting and then joining is not allowed – a bucket which was created by a split cannot be joined with another bucket in the same epoch.
Handling new minima and maxima.
If the buffer contains a new minimum $\tau^{\prime}_{1}$ (an item smaller than the previous minimum $\tau_{1}$), we add a new bucket $(-\infty,\tau^{\prime}_{1}]$ before all of the existing buckets, set its counter to the frequency of $\tau^{\prime}_{1}$ in the buffer. We also increase the counter of the bucket $(\tau^{\prime}_{1},\tau_{1}]$ by the number of buffer items in this interval. A new maximum $\tau^{\prime}_{k}$ is handled similarly, by adding a new bucket $(\tau_{k},\tau^{\prime}_{k}]$ and setting its counter to the number of buffer items in this interval. For each of these at most two new buckets, we perform a join of another pair, selected according to the rules described below. Note that these at most two new buckets may be later split during Consolidate in which they were created.
Maintaining bounds on bucket counters and small heuristic error.
We maintain two invariants, the first of which is used to prove worst-case bounds while the second ensures high accuracy in practice.
In order to obtain a low worst-case uniform error, we need to ensure that no bucket has a too large counter. To this end, we require that all bucket counters are bounded by $O(n/k)$, specifically
(1) $\forall i=1,\dots,k:\,b_{i}\leq C_{b}\cdot n/k$
for some constant $C_{b}>1$ that will be determined in the analysis in App. B.3.
Secondly, we define a heuristic error for each bucket, which is intended to capture “hard” parts of the input, where the input distribution changes a lot, namely, where the second derivative of the empirical CDF is large. For such buckets, the PCHIP interpolation may have larger error than for buckets with a low second derivative. Since the empirical PDF (i.e., first derivative of the CDF) in the bucket $i$ is $b_{i}/\ell_{i}$, we use the following heuristic error for bucket $i>1$, which approximates the second derivative of the empirical CDF, normalized by length squared:
(2) $\max\left(\frac{\left|\frac{b_{i}}{\ell_{i}}-\frac{b_{i-1}}{\ell_{i-1}}\right|}{\ell_{i}+\ell_{i-1}}\,,\,\frac{\left|\frac{b_{i+1}}{\ell_{i+1}}-\frac{b_{i}}{\ell_{i}}\right|}{\ell_{i+1}+\ell_{i}}\right)\cdot\ell_{i}^{2}.$
For the edge cases $i=2$ and $i=k$, we make the following assumptions: $b_{k+1}=0$, $\ell_{1}=\ell_{2}$, and $\ell_{k+1}=\ell_{k}$; that is, we add a “virtual” buckets $k+1$ with zero counter and a length equal to the length of the adjacent (real) bucket, and we set the length of bucket 1 to the length of bucket 2.
We now describe how we use the bucket bound (1) and heuristic error (2) when performing the splits and joins. Each time we split a bucket $i$, we choose a pair of adjacent buckets that do not contain buckets created from splitting $i$ and we join that chosen pair in order to keep the total number of buckets the same. The pair is selected so that the bucket resulting from joining has the lowest heuristic error of all joinable pairs, defined as follows. In a nutshell, the bucket resulting from the join should not violate the bucket size bound (1) or be relatively close to it. Moreover, none of the two buckets should be a subject to a split recently, which we implement using the protected thresholds.
Definition 1.
We say that a pair $i,i+1$ of adjacent buckets is joinable if both of the following conditions hold:
  1. (i)
    the counter of the bucket resulting from the join satisfies $b_{i}+b_{i+1}\leq 0.75\cdot C_{b}\cdot n/k$, and
  2. (ii)
    threshold $\tau_{i}$ is not protected.
We perform splits as follows. First, we take all buckets that exceed the size bound (1). We split them one by one, always performing a join satisfying conditions of Definition 1 for each split; in Lemma B.2 we show that at least one pair of adjacent buckets can be joined. We also join at most two further pairs of joinable buckets if the buffer contains a new minimum or maximum, each creating a new bucket. Note that a bucket may need to be split repeatedly during one call of Consolidate, namely, if it exceeds the bucket bound (1) substantially.
After all these necessary splits and joins, we consider a bucket $i$ with a high heuristic error (2), and we split it if its error is more than $\gamma$ times greater than the lowest heuristic error of a bucket resulting from joining a joinable pair, for a parameter $\gamma>1$ (say, $\gamma=1.5$). Furthermore, we do not perform a split of a bucket $i$ due to the heuristic error if there are less than $k/3+2$ pairs of adjacent buckets that can be joined, which is needed to prove that there is a joinable pair of buckets; see Lemma B.2. Finally, if the counter of bucket $i$ is substantially below the bound, say, $b_{i}\leq(C_{b}/100)\cdot n/k$, we also do not split bucket $i$, avoiding splitting almost empty buckets. We summarize the conditions for splitting:
Definition 2.
We say that a bucket $i$ is splittable if one of the following conditions holds:
  1. (i)
    bucket $i$ violates the bucket bound (1), i.e., $b_{i}>C_{b}\cdot n/k$, or
  2. (ii)
    there are at least $k/3+2$ joinable pairs of buckets, $b_{i}>(C_{b}/100)\cdot n/k$, and the heuristic error (2) of bucket $i$ is $\gamma>1$ times larger than the heuristic error of a joinable pair of buckets (after joining) not overlapping bucket $i$.
We repeat finding a splittable bucket $i$ and a joinable pair of buckets not overlapping bucket $i$, performing the split and join as long as there are splittable buckets. Since every split increases the number of protected thresholds, we know that this process terminates after at most $O(k)$ iterations. In Appendix B.2, we show it is also well-defined.
Lemma 3 (Lemma 2).
SplineSketch always contains a joinable pair of adjacent buckets, according to Definition 1.
Input : Buffer of size $n$, parameter $k$ (number of buckets)
Output : Initialized $k$ buckets with thresholds $\tau_{1},\dots,\tau_{k}$ and counters $b_{1},\dots,b_{k}$
1 Sort all $n$ items in the buffer in ascending order;
2 For $i=0,\dots,k-1$, pick the item at index $\min\{\lceil i\cdot(n-1)/(k-1)\rfloor,n-1\}$ from the sorted buffer to form $k$ thresholds $\tau_{1},\dots,\tau_{k}$;
3 For each bucket $(\tau_{i-1},\tau_{i}]$, set the counter $b_{i}$ to the exact number of items in the buffer that fall into this interval;
4 Return the initialized buckets with thresholds $\tau_{1},\dots,\tau_{k}$ and counters $b_{1},\dots,b_{k}$;
Algorithm 1 Buckets Initialization
Input : $k$ buckets and buffer of new items
Output : Updated buckets and protection bit vector
1
2if $n\geq t$ then
3   Reset protection bit vector, and set $t:=1.25\cdot t$;
4
5Incorporate buffer items into each bucket’s counter, and empty the buffer;
6 if new minimum or maximum found in the buffer then
7   Add a new boundary bucket for minimum/maximum;
8   For each new bucket, find a pair of joinable buckets $j,j+1$, and join this pair;
9
10while exists a splittable bucket $(\tau_{i-1},\tau_{i}]$, first processing those exceeding the bucket bound (1) do
11   Split bucket $i$ into $(\tau_{i-1},\tau^{\prime}]$ and $(\tau^{\prime},\tau_{i}]$ for $\tau^{\prime}=(\tau_{i}+\tau_{i-1})/2$;
12   Mark thresholds $\tau_{i-1},\tau^{\prime},\tau_{i}$ as protected;
13   Join a pair of joinable buckets $j,j+1$ (Def. 1) that does not overlap with $(\tau_{i-1},\tau^{\prime}]$ and $(\tau^{\prime},\tau_{i}]$ and has the lowest heuristic error;
14
Algorithm 2 Consolidate
Resizing the sketch during Consolidate.
Bucket consolidation allows to change the number $k$ of buckets, making the sketch larger or smaller as desired. Changing from $k$ to $k^{\prime}$ buckets is straightforward: If we increase the sketch size, i.e., $k^{\prime}>k$, we perform $k^{\prime}-k$ splits without any join, first dealing with buckets exceeding bound (1) and then in the order of the heuristic error. On the other hand, if $k^{\prime}<k$, after performing the necessary splits due to exceeding the bound (and the corresponding joins), we execute $k-k^{\prime}$ joins in the order of their heuristic errors after joining, on bucket pairs satisfying Definition 1 and without performing any splits. These separate splits or joins may be followed by performing Consolidate in a normal way, where we always join after splitting. If $k^{\prime}$ is substantially different from $k$, say $|k^{\prime}-k|>0.25\cdot k$, we also reset the protection bit vector.
3.2. Merging Two SplineSketches
We describe how to merge two SplineSketches $S_{1}$ and $S_{2}$ with $k_{1}$ and $k_{2}$ buckets, respectively, into one sketch $S$ for the combined datasets. We merge one sketch into the other, namely $S_{2}$ into $S_{1}$, but the user can specify whether to merge $S_{2}$ into $S_{1}$, or vice versa. The default choice is that we merge the sketch summarizing a smaller number of items into the one with more items. Furthermore, in the analysis, we assume that $k_{1}=k_{2}$ but the merge operation works even for $k_{1}\neq k_{2}$.
First, we create a buffer that contains all the items from the buffers of the two sketches we are merging. Second, we take the union of all the thresholds $\tau$ from both sketches. After removing duplicates, this gives up to $k_{1}+k_{2}$ thresholds. For each of these thresholds $\tau$, we query for its rank in both input sketches and add these two ranks to get the rank of $\tau$ in the merged sketch $S$. We set the counter of each bucket to be the difference between the ranks of the endpoints. Third, the new sketch inherits the bit vector for threshold protection from $S_{1}$, meaning that any threshold taken from $S_{1}$ remains protected if it was protected in $S$. However, if the total input size of the new sketch is larger than the epoch end for $S_{1}$, we set all thresholds to the unprotected state, starting a new epoch.
Finally, we run Consolidate that reduces the number of buckets to $k_{1}$ and also merges the buffer to buckets. Namely, we repeatedly take a joinable pair of adjacent buckets whose joining would result in a bucket with the lowest heuristic error and join it without any split until we have $k_{1}$ buckets.
3.3. Theoretical Properties
We now provide error bounds that hold for any input in the mergeability setting, which captures the streaming model as we can interpret stream updates as merges with a single-item sketch.
Theorem 4 (See Appendix B.3).
Suppose that we build a SplineSketch using any sequence of pairwise merge operations executed on $n$ data items, with every intermediate SplineSketch consisting of $k$ buckets. Then the resulting SplineSketch has a worst-case rank error of $O(\log(\alpha)\cdot n/k)$.
The analysis assumes that no element appears more than $n/k$ times. In order to lift this assumption, we can run a heavy-hitter Misra-Gries (MG) sketch (Misra and Gries, 1982) of size $O(k)$ in parallel. We describe how to effectively combine it with SplineSketch in Sec. 4.2.
We note that state-of-the-art quantile sketches achieve better worst-case accuracy-space trade-offs: Namely, KLL using space of $O(k)$ memory words has worst-case rank error $O(n/k)$ with constant probability (Karnin et al., 2016). The deterministic Greenwald-Khanna (GK) sketch guarantees rank error $O(n/k)$ but with worst-case space $O(k\cdot\log(n/k))$ (Greenwald and Khanna, 2001); the $\log(n/k)$ factor is in general incomparable to the $\log(\alpha)$ in the error for SplineSketch. However, both KLL and GK achieve significantly worse error in practice, as we show in Sec. 5. As for mergeability, KLL is fully mergeable while retaining the guarantees (Karnin et al., 2016) but GK is known to be only one-way mergeable (Agarwal et al., 2013). Other quantile sketches have worse rank error or do not provide a bounded rank error; see Sec. 2.
Our analysis is very flexible with respect to individual components and parameters of the sketch, with only constant factor changes (such as the value of $C_{b}$ in Eqn. 1). Namely, it is independent of how the heuristic error is computed (and of the value of $\gamma$ in Def. 2) and as for other parameters, the analysis needs that epoch lengths increase geometrically. As for the interpolation, it can be changed to linear or other monotone interpolation that is exact on the thresholds if it satisfies that in the middle of a non-empty bucket, the interpolation is bounded away from the minimum and maximum estimated values inside the bucket, i.e., a bucket counter is divided somewhat evenly during a split (a property similar to Lemma 1 in Appendix A).
Regarding the effect of resizing the sketch from $k^{\prime}$ buckets to $k^{\prime\prime}$ buckets on the guarantees, the error bound of $O(\log(\alpha)\cdot n/k)$ holds with $k=\min\{k^{\prime},k^{\prime\prime}\}$ after performing the resizing. To the best of our knowledge, resizability is not addressed in the literature on quantile summaries yet.
As for the time complexities, the time for one rank or quantile query is $O(\log k)$, as derived at the beginning of Sec. 3. The amortized update time can be made $O(\log k)$ by a heap-based implementation of Consolidate, as explained in Sec. 4.1. Finally, the merge time is $O(k\cdot\log k)$, again for the heap-based Consolidate. These theoretical time complexities are asymptotically the same as for KLL, GK, and $t$-digest, albeit particular implementations of the sketches may not satisfy the theoretical bounds.
4. Implementation Details
Here, we explain the details of our SplineSketch implementation (in Python and Java) and elaborate on the update time.
First, we naturally use the double data type (64-bit floating point numbers) for thresholds. This, however, implies some limitations, such as worse accuracy for some inputs with many significant digits (e.g., the SOSD benchmark (Kipf et al., 2019; Marcus et al., 2020)), which is due to the fact that many different items get rounded to the same double value. This can be addressed by changing the underlying data type for thresholds, e.g., to 128-bit floating-point numbers or to unsigned 64-bit integers. Note that the internal PCHIP interpolation would still need to use floating-point numbers.
Quantile queries.
So far, we focused on answering rank queries. In order to implement a quantile query, i.e., returning an estimated $r$-th smallest number for a given $r$, we invert the interpolation, and query the inverse function at $r$; the inverse exists since buckets are non-empty and the PCHIP interpolation is thus strictly increasing from $\tau_{1}$ to $\tau_{k}$. The inverse function can be numerically evaluated by using, for example, the Newton’s method or even just binary search. Note that our algorithm, similarly as $q$-digest (Shrivastava et al., 2004), may return a number that did not appear in the stream, unlike comparison-based methods such as the KLL sketch (Karnin et al., 2016) or the Greenwald-Khanna sketch (Greenwald and Khanna, 2001).
4.1. Implementations of Bucket Consolidation
We describe two particular implementations of Consolidate. The first achieves a low amortized update time, while the second is easier to implement and works well in practice.
Heap-based Consolidate.
To obtain an asymptotically fast Consolidate, one can use a combination of doubly-linked lists for maintaining buckets and two heaps, sorted by the bucket size and heuristic error, respectively. This approach is formalized in the following lemma. Since the buffer size is $O(k)$, we get $O(\log k)$ amortized complexity of an update with heap-based Consolidate.
{restatable}
lemmaheapBasedConsolidate Consolidate can be implemented in $O(k\log k)$ time.
Proof.
We maintain buckets in a doubly linked list $L$. We use a queue $Q_{S}$ to keep track of buckets that need to be split due to violating bucket bound. We also use two heaps: a max-heap $H_{S}$ to keep track of buckets that might get split due to the heuristic error (2), ordered by their heuristic errors, and a min-heap $H_{J}$ for joinable pairs of buckets, ordered by their heuristic errors after temporarily joining each pair. In both heaps, every node contains a pointer to the corresponding bucket (or pair of buckets) in $L$. Conversely, every bucket in $L$ contains a (constant size) list of pointers to nodes of $H_{S}$ and $H_{J}$ containing that bucket. Both heaps can be initialized in $O(k)$ time.
After every split and join performed by Consolidate, we update $Q_{S},H_{S}$, and $H_{J}$ according to the new bucket sizes. Since every split and join affects only the sizes and heuristic errors of a constant number of adjacent buckets, the update can be done in $O(1)$ time using the pointers that we maintain. The modifications of the keys and deletions of arbitrary nodes in the heaps can be done in $O(\log k)$ time using standard decrease/increase key operation.
Every split introduces a constant number of new protected thresholds and every join is accompanied with a corresponding split. Since the number of possible protected thresholds is bounded by $O(k)$, the total number of splits and joins performed by Consolidate is also $O(k)$. Therefore, the total time complexity of the heap based implementation of Consolidate is $O(k\log k)$. ∎
Iteration-based Consolidate.
In our implementation, we use a practically more efficient Consolidate without the need for linked lists and heaps. We process the buckets in iterations such that each iteration splits or joins one bucket at most once and not both at the same time; note that more iterations are possibly needed as one bucket may be split or joined multiple times during Consolidate. We stop when the last iteration does not change any threshold. Throughout this process, we keep the buffer and the old buckets for more accurate setting of counters for the new buckets.
In each iteration, we first collect the buckets with size exceeding the bound (1) into a set $E$. We also keep a list of splittable buckets (Def. 2), sorted non-increasingly by their heuristic error, and a list of joinable pairs of buckets (Def. 1), sorted non-decreasingly by their heuristic errors after temporarily performing their joins. We select $|E|$ joinable pairs according to their order that do not overlap buckets in $E$ or a previously selected joinable pairs. Next, from the remaining buckets (not yet split or joined), we consider the splittable bucket with the largest heuristic error $h_{\mathrm{split}}$ (if any) and the joinable pair with the lowest heuristic error $h_{\mathrm{join}}$ and if $h_{\mathrm{split}}>\gamma\cdot h_{\mathrm{join}}$ for $\gamma>1$, then we perform the split and join of these buckets; in our implementation, we use $\gamma=1.5$. After collecting non-overlapping sets of buckets to split and join, we perform these splits and joins in one pass over the buckets, using the buffer and interpolation over the buckets in their state before executing Consolidate to estimate ranks in the middle of split buckets; the usage of former buckets is important to decrease the error accumulation. We refer to the prototypes referenced in Sec. 5 for details.
Finally, we remark that the multiplicative constant $C_{b}$ in the bucket bound (1) resulting from the analysis in App. B is too large for practical usage, and we just use $C_{b}=3$, while increasing $C_{b}$ during an epoch if needed and resetting it to $C_{b}=3$ at each epoch end. That is, if there are no joinable pairs when a bucket must be split due to exceeding the bound, we double $C_{b}$ and reset the iteration; we note that such increases of $C_{b}$ happen rarely in practice.
4.2. Handling Frequent Items
The theoretical error guarantees of SplineSketch, outlined in Sec. 3.3, require a method for filtering items with frequency above $n/k$. This can be done using a heavy hitter sketch such as the Misra-Gries (MG) summary (Misra and Gries, 1982). It is also important in practice, as our experiments in Sec. 5.3 demonstrate. However, using the MG sketch comes at a cost of increased update, merge, and query times, and therefore, we provide two versions of SplineSketch, with and without MG.
SplineSketch with MG
The MG sketch stores up to k1k-1 items, each with a counter $c_{x}$. When a new element $x$ arrives, if $x$ is already in the sketch, its counter is incremented. Otherwise, if $x$ is not in the sketch, we add it with a counter of $1$. If adding a new item causes the sketch to reach its capacity of $k$ items, all counters are decremented. Any item with a counter that hits zero is then removed. It can be shown that the counters in the sketch underestimate the true frequencies by at most $n/k$. In particular, any element with a frequency greater than $n/k$ is present in the sketch (Misra and Gries, 1982). The MG sketch is also fully mergeable (Agarwal et al., 2013).
In our implementation, each item $x$ in the MG sketch has an additional counter, $\hat{c}_{x}$, which is never decreased. It represents the number of times $x$ has been inserted since the beginning of the stream or its last removal from the MG sketch. During Consolidate, the MG sketch processes a buffer of new items in two stages. First, we filter out buffer items that are present in MG and increase their sketch counters by the buffer frequencies. For the remaining distinct items in the buffer, we process them one by one, and add them to the sketch with both their counters ($c_{x}$ and $\hat{c}_{x}$) set to their buffer frequency. If at any point the MG sketch becomes full, we decrease the $c_{x}$ counters of all items in the MG sketch by the smallest counter. We then remove all items with $c_{x}=0$ and add them back into the buffer with frequency $\hat{c}_{x}$. Finally, the items that remain in the buffer are processed by standard Consolidate.
Last, after processing the whole input, the extra space used by the MG sketch is compressed by removing all the items $x$ with frequency $\hat{c}_{x}<n/(2\cdot k)$. Additionally, if $\ell$ is the number of remaining items with $\hat{c}_{x}\geq n/(2\cdot k)$, we resize SplineSketch so that there are $\max\{k-\ell,k/2,6\}$ buckets. Thus, if $\ell\leq k/2$, there will be $k$ buckets and MG items in total. Based on the standard analysis of the Misra-Gries algorithm (Misra and Gries, 1982), each input element is inserted into the SplineSketch at most $n/k$ times. To process a rank query $y$ in the MG sketch, we sum the frequencies $\hat{c}_{x}$ for all items $x\leq y$. This sum is not an approximation with respect to the items summarized by the MG sketch, so it does not increase the error. We then add the result of the SplineSketch query for $y$, i.e., the $y$’s rank estimate based on the buckets and the buffer.
SplineSketch without MG
Since item frequencies can be relatively small in many applications, we also implement a faster and simpler version without heavy-hitter filtering. Still, this version tries to apply a heuristic detection of buckets with high-frequency items.
Namely, we set a lower limit on the relative length of a bucket, depending on numerical precision $\delta$ of data. That is, a bucket $(\tau_{i},\tau_{i+1}]$ must have length $\tau_{i+1}-\tau_{i}\geq\delta^{\prime}\cdot\max\{|\tau_{i}|,|\tau_{i+1}|,\zeta\}$, where $\delta^{\prime}>\delta$ is a parameter and $\zeta>0$ is the smallest non-zero absolute value of an input item ($\zeta$ is intended to avoid too small buckets around 0). In our prototype implementations with the double data type, we set $\delta^{\prime}=10^{-8}$. We do not split a bucket that would violate this relative length bound, even if it does not satisfy the size bound (1); this way, we avoid too many splits of a bucket with a frequent item.
We also deal with repeated thresholds when creating initial buckets in Alg. 1, by removing thresholds that would violate the bucket length bound, replacing them by new thresholds relatively close to the previous threshold so that the resulting buckets would have length close to the relative length bound and we end up with $k$ thresholds. This in fact heuristically guarantees that a frequent item gets its own bucket if it is frequent from the beginning.
5. Experimental Evaluation
We have implemented SplineSketch as a prototype in Python and Java and compared the Java version with state-of-the-art quantile sketches for the uniform rank error with an available implementation, namely, $t$-digest (Dunning and Ertl, 2019), MomentSketch (Gan et al., 2018), and KLL (Karnin et al., 2016). Our implementation of SplineSketch and the code to run the experiments or generate the synthetic datasets, is available at supplementary repository (https://github.com/PavelVesely/SplineSketch-experiments). We measure the average and maximum absolute rank errors for $10^{5}$ evenly spaced queries, together with average update (or merge if applicable) and query times. Experiments were performed on an AMD EPYC 7302 (3 GHz) server with 251 GB RAM and SSD disk, and run in memory, with the datasets, queries, and sketches’ output stored on disk.
Quantile sketches.
We have used the Java prototype of SplineSketch, in two versions: with and without the Misra-Gries sketch. We have also evaluated $t$-digest (Dunning and Ertl, 2019) (v3.3, https://github.com/tdunning/t-digest), MomentSketch (Gan et al., 2018) (https://github.com/stanford-futuredata/msketch), and KLL (Karnin et al., 2016) (Java implementation by Apache DataSketches, v6.0.0, https://github.com/apache/datasketches-java), and GKAdaptive, a practically well-performing variant of the Greenwald-Khanna sketch (Greenwald and Khanna, 2001), implemented in Java according to (Luo et al., 2016). For $t$-digest, we use the default merging variant with $k_{0}$ scale function that aims at the uniform error (Dunning and Ertl, 2019). KLL was used in the variant with the double data type, i.e., an instance of KllDoublesSketch. We note that the MomentSketch implementation only answers quantile queries, while our experiments required rank estimates; we have simulated rank queries using binary search that was done for all $m$ queries in parallel in order to decrease the number of calls to the quantile query method from $\approx m\log_{2}n$ (querying just one rank) to $\approx\log_{2}n$ calls (querying $m$ ranks), to avoid repeated computation of the distribution fitting the moments and log-moments, which is a costly operation. We have not included other quantile sketches because they have worse error than KLL or t-digest, lack available implementations (e.g., algorithms from (Schiefer et al., 2023; Gribelyuk et al., 2025)), or aim for different error guarantees; namely, ReqSketch (Cormode et al., 2023) provides relative rank error and its space usage is much worse than KLL’s, while the algorithmic technique is similar. DDSketch (Masson et al., 2019) and UDDSketch (Epicoco et al., 2020) aim for the relative value error, which captures the distribution tails but unlike the rank error, fails to provide a reasonable approximation for most of the distribution; indeed, we have verified that the uniform rank error of DDSketch is 10-100 times worse than KLL and even more compared to SplineSketch; see the supplementary repository for details.
The sketch size is measured in bytes when serialized on disk for storage without supporting data structures such as buffers. Namely, for SplineSketch and $t$-digest we count 16 bytes per bucket/centroid, for MomentSketch we measured $16k+16$ bytes ($k$ moments and log-moments, minimum, and maximum), for KLL we count 8 bytes per stored item, and for GKAdaptive, we count 24 bytes per a stored tuple of one item and its rank lower and upper bounds. For SplineSketch with MG, we also account for the MG sketch size, after removing non-frequent items from the MG sketch and reducing the number of buckets based on the number of heavy hitters as described in Sec. 4.2.
Datasets.
We use three real-world datasets, the first two from UCI Machine Learning Repository (Kelly et al., 2023) : the HEPMASS dataset (Whiteson, 2016) (n=10 500 000n=10\,500\,000 items from the 2nd column), the Power dataset (Hebrail and Berard, 2006) (n=2 075 259n=2\,075\,259 items from the 3rd column), and the Books dataset from SOSD (Kipf et al., 2019; Marcus et al., 2020). We have synthetically generated 7 datasets by drawing i.i.d. (independent and identically distributed) samples from a range of distributions: uniform, normal, Pareto, Gumbel, lognormal, loguniform, and randomly signed loguniform. We have also generated three datasets with distribution change, all starting with $n/2$ items from a normal distribution and then either changing the parameters of the normal distribution (two options for the parameter change) or adding samples from a set of 42 distinct items, which will then all have high frequency. Finally, we generated a sorted input with frequent items, each with with random frequency between 1 and $n/50$. All synthetic datasets for accuracy experiments consist of $n=10^{8}$ items.
Rank queries are generated by equally spaced selection from sorted data and are executed in one batch whenever possible. While all of the sketches allow any number to be queried, using only the data points for queries leads to the same results.
5.1. Evaluation in the Streaming Setting
Accuracy-space trade-off.
We first present the results in the streaming setting, focusing on synthetic and real-world inputs without frequent items. Figs. 1 and 4 show the average errors of the sketches depending on their size in bytes. Fig. 5 shows the maximum error over all queries, again depending on the sketch size. In all these plots, the error is normalized by the input size, i.e., it is the difference between the true rank and the estimated rank of a query item divided by the input size $n$, averaged or maximized over $10^{5}$ queries. We also depict the theoretically optimal error bound of $O(1/k)$ in purple (loosely dotted), where $k$ is the sketch size and the hidden constant factor is based on the maximum error of KLL and GKAdaptive.
Overall, SplineSketch consistently provides the best accuracy if given sufficient size, namely $k\geq 100$, even though in rare cases or for small sketch sizes, $t$-digest or MomentSketch are better. In more detail, compared to $t$-digest, SplineSketch typically achieves 2-20 times smaller error, and in some cases, such as for normally distributed data, up to 100 times. In a few cases, their error was similar, e.g., the maximum error for signed loguniform distribution with large exponents. We note that the final $t$-digest size is somewhat unpredictable as the final number of centroids does not match the compression parameter given to it, sometimes by 66%. This demonstrates that $t$-digest does not use the space budget efficiently, unlike SplineSketch which always has $k$ buckets.
MomentSketch is mainly intended for use with very small size, typically up to $k=15$ moments and log-moments, when it often gives the best accuracy by 1-2 orders of magnitude. It works the best on many smooth distributions (uniform, normal, loguniform, etc.). However, its performance is significantly affected by non-smoothness as witnessed on the real-world datasets. Moreover, one cannot increase accuracy by increasing $k$ because of numeric issues. Furthermore, MomentSketch query procedure failed with an exception on many datasets, specifically on signed loguniform distribution with large exponents for any $k$ and on other datasets for large $k$. Finally, the KLL sketch and GKAdaptive are worse than SplineSketch typically by 1-3 orders of magnitude as they use no interpolation (cf. (Schiefer et al., 2023) for KLL with interpolation; however, no implementation is available). Nevertheless, as KLL and GKAdaptive are comparison-based, their errors depend just on the data size and the arrival order, not on their particular distribution, and their error decreases linearly.
Finally, we note that both versions of SplineSketch (without and with the MG sketch) provide the same accuracy on the inputs in Figs. 4 and 5.
Update and query times.
[…]
(a) Normal distribution
(b) Uniform distribution
(c) Pareto distribution
(d) Lognormal distribution
(e) Loguniform distribution
(f) Signed loguniform distribution with large exponents
(g) Normal distribution with a small parameter change after $n/2$ items
(h) Normal distribution with a large parameter change after $n/2$ items
Figure 4. Average rank error (log-scale) depending on the sketch size in bytes (average taken over $10^{5}$ evenly spaced queries).
(a) HEPMASS dataset (Whiteson, 2016)
(b) Power dataset (Hebrail and Berard, 2006)
(c) Books dataset (Kipf et al., 2019; Marcus et al., 2020)
(d) Normal distribution
(e) Uniform distribution
(f) Pareto distribution
(g) Lognormal distribution
(h) Loguniform distribution
(i) Signed loguniform distribution
(j) Signed loguniform distribution with large exponents
(k) Normal distribution with a small parameter change after $n/2$ items
(l) Normal distribution with a large parameter change after $n/2$ items
Figure 5. Maximum rank error (log-scale) depending on the sketch size in bytes, on a selection of datasets. The maximum is taken over $10^{5}$ evenly spaced queries.
[…]
(a) Log-log plot with time per query depending on the number of queries.
(b) Time per update depending on sketch size.
(c) Time per query depending on sketch size for $10^{5}$ queries.
[…]
5.2. Evaluation of Mergeability
We evaluated the performance sketches when they are created by pairwise merge operations. To this end, we split the input into 10 000 chucks, feed each chunk into one sketch, and then merge the resulting sketches into one in a binary tree fashion, i.e., merging sketches summarizing a similar number of items.
[…]
Figure 8. Average rank error (log-scale) depending on the sketch size in bytes, on the normal distribution for sketches built by merge operations from sketches for 10 000 input chunks.
Figure 9. Time per merge operation, in the same setting as in Figure 8.
5.3. Inputs with Heavy Hitters
Here, we focus on inputs with high-frequency items, where SplineSketch without the MG sketch does not perform so well. Our results in Figs. 10(a) and 10(b) confirm that using the Misra-Gries (MG) sketch alongside SplineSketch does indeed restore the very high accuracy when the MG sketch is large enough to detect the heavy hitters (note that the size of the MG sketch after compressing as described in Sec. 4.2 is accounted for). In more detail, while the error of $t$-digest and SplineSketch without MG gets worse and does not improve with the sketch size, SplineSketch with MG achieves by up to three orders of magnitude better error, and zero error on inputs consisting of a small number of distinct items only as in Fig. 10(b). We note that the error of KLL and GKAdaptive remains the same as on inputs with distinct items as they are comparison-based. As shown in Fig. 10(c), the update time of SplineSketch with MG is about three times worse than without MG (though it can be optimized), while merge and query times are only slightly larger.
(a) Average rank error (log-scale), on normal distribution with $n/2$ items, followed by $n/2$ high-frequency items.
(b) Average rank error (log-scale), on a sorted input with frequent items.
[…]
Figure 10. Accuracy and update time for inputs with frequent items, depending on the sketch size in bytes.
5.4. Ablation Studies
Here, we provide insights into the specific design choices in SplineSketch. First, we evaluated the benefit of using PCHIP interpolation and found that PCHIP interpolation achieves 3–10 times better error than just linear interpolation over the buckets; see Fig. 11(a).
Second, we focused on the heuristic error and compared several variants instead of the second derivative of the empirical CDF, normalized by length squared, defined in (2). The evaluated alternatives are: using no heuristic error, bucket length, bucket counter (which is the first derivative of the empirical CDF), and the third derivative of the empirical CDF, normalized by length cubed. Fig. 11(b) shows that (2) is indeed the best, while using the bucket length, i.e., making the buckets similar in length on the input range, is only about 20% worse. The third derivative is about 1.5–2 times worse, and using the bucket counter or no heuristic error is over 10 times worse.
Finally, we briefly comment on parameter tuning. We did not observe significant differences in the accuracy of SplineSketch with respect to parameter $\gamma>1$ in Definition 2 (for values between 1.1 and 2). For the initial multiplicative constant $C_{b}$ in the bucket bound (1), we tested values in $[1,7]$ and for $C_{b}\geq 3$, the accuracy was similar ($C_{b}<3$ was only better on the uniform distribution). Other parameters, such as the epoch increase factor (default=1.25) or the minimum fraction of the bucket bound (default=0.01) also do not influence the accuracy much if adjusted. See the supplementary repository for details.
(a) Interpolation.
(b) Heuristic error.
Figure 11. Ablation studies on the normal distribution.
6. Conclusions and Discussion
In this work, we designed a new deterministic quantile sketch which has fixed memory consumption $O(k)$. The main technical contribution is in careful maintenance of bucket thresholds and counters in order to obtain extremely accurate rank and quantile estimates, with uniformly bounded worst-case error, achieving a near-optimal accuracy-space trade-off. We have proven that our sketch is fully mergeable, rendering it suitable for parallel or distributed processing of massive datasets.
Besides very high accuracy on various datasets, the key advantage of our approach is flexibility. For instance, it is possible to resize the sketch, making it smaller or larger depending on how the memory resources change (e.g., one can use larger size $k$ for input processing than for storage). More importantly, if the user has a prior knowledge about the data distribution, it is possible to preset the initial thresholds of SplineSketch based on this prior distribution and aim for even lower error. Furthermore, if the prior knowledge turns out to be wrong, our sketch adapts to the new distribution, similarly as we have shown in Fig. 4(h).
One of the challenges when implementing SplineSketch is dealing with high-frequency items. One possible approach is using the Misra-Gries sketch (Misra and Gries, 1982), which restores theoretical guarantees (Sec. 3.3) and very high accuracy for inputs with heavy hitters, as demonstrated in Section 5.3. Nevertheless, using a heavy-hitter sketch comes at a cost of increased update, merge, and query times, so we also provide a faster version without a heavy hitter sketch. Another limitation of our implementation arises from using the double data type in Java, which means that for some inputs with many significant digits (e.g., the SOSD benchmark (Kipf et al., 2019; Marcus et al., 2020)), accuracy degrades as the sketch rounds many items to the same double value; this can be addressed by changing the underlying data type, e.g., to 64-bit integers. While we optimized the Java version to some extent, our implementation is so far a prototype intended to demonstrate the main advantages and challenges of our approach. Due to the intricacies of maintaining the buckets, our sketch is slower in processing the input stream than KLL or MomentSketch but already faster in streaming than $t$-digest. However, further optimizations are possible, such as the ones presented in (Pfeil et al., 2025) (though not applicable to Python or Java). Thus, a well-optimized implementation may be significantly faster than $t$-digest.
From the theoretical point of view, we ask whether one can modify our algorithm so that the error bounds have better dependency on $\alpha$. We note that there is some evidence from the comparison-based lower bounds in (Cormode and Veselý, 2020) that a factor of $\log\log\alpha$ may be necessary without using bit-packing tricks developed for $q$-digest in (Gupta et al., 2024). The accuracy in practice, despite being already very high, may also be improved, and it is not clear where is the accuracy limit of quantile summaries on real-world datasets.
Our work opens up a new direction for future work in quantile streaming estimation. In particular, we believe that our techniques have potential to yield much better quantile sketches with relative-error guarantees, capturing the distribution tails more accurately than the median. However, achieving the relative error is substantially harder as witnessed by several major open problems in the theory of relative-error quantile sketches (cf. (Cormode and Veselý, 2020; Gribelyuk et al., 2025)).
Acknowledgments.
We thank the reviewers for their constructive feedback and suggestions for improvement. ChatGPT was used to generate parts of this work, including polishing language in the text, pseudocodes prototyping, and translation of SplineSketch from Python to Java. We have checked all the LLM generated content, corrected it where necessary. In particular, the Java implementation was largely rewritten and optimized manually.
Appendix A Bounds on Middle Point of Monotonic Cubic Polynomials
The following lemma provides a key technical property of PCHIP interpolation that we use subsequently in the analysis of our algorithm in Appendix B.
Lemma 1.
Let $P(x)$ be a cubic polynomial such that $P(0)=0$, $P(1)=1$, and $P(x)$ is non-decreasing on the interval $[0,1]$. Then, $P\left(\dfrac{1}{2}\right)\in[\beta,1-\beta]\approx[0.066,0.933],$ where $\beta=\frac{1}{2}-\frac{\sqrt{3}}{4}\approx 0.067$.
Proof.
We prove that $P(\dfrac{1}{2})\leq\frac{1}{2}+\frac{\sqrt{3}}{4}$. The other inequality then holds by symmetry, that is, considering the polynomial Q(x)=1P(1x)Q(x)=1-P(1-x) which is also non-decreasing and $Q(\dfrac{1}{2})=1-P(\dfrac{1}{2})$. From now on, we focus on proving that $P(\dfrac{1}{2})\leq\frac{1}{2}+\frac{\sqrt{3}}{4}$.
Let us denote the coefficients of $P$ by $a,b,c$ and $d$, that is,
$P(\lambda)=a\lambda^{3}+b\lambda^{2}+c\lambda+d\,.$
Because $P(0)=0$, we have that $d=0$. At the same time, because $P(1)=1$, we have that $a+b+c=1$. From the assumption that the polynomial is non-decreasing, we also have that its derivative is non-negative, meaning that for any $\lambda\in[0,1]$,
$3a\lambda^{2}+2b\lambda+c\geq 0\,.$
At $1/2$, the polynomial evaluates to
f(a,b,c)=a/8+b/4+c/2.f(a,b,c)=a/8+b/4+c/2\,.
We need to prove that under the constraints we mentioned, it holds that
$f(a,b,c)\leq\frac{1}{2}+\frac{\sqrt{3}}{4}$
First, using the constraint $a+b+c=1$, we express $c$ as c=1abc=1-a-b. Substituting this into the inequality $3a\lambda^{2}+2b\lambda+c\geq 0$, we obtain $3a\lambda^{2}+2b\lambda+(1-a-b)\geq 0$. Simplifying this expression gives:
(3) $(3\lambda^{2}-1)a+(2\lambda-1)b+1\geq 0\,.$
Next, we choose $\lambda=\dfrac{3+\sqrt{3}}{6}$, which lies within $[0,1]$. We compute the coefficients:
$3\lambda^{2}-1=\dfrac{\sqrt{3}}{2},\quad 2\lambda-1=\dfrac{\sqrt{3}}{3}\,.$
Substituting these values into inequality (3) gives:
$\dfrac{\sqrt{3}}{2}a+\dfrac{\sqrt{3}}{3}b+1\geq 0\,$,
which simplifies to:
(4) $\sqrt{3}\left(\dfrac{a}{2}+\dfrac{b}{3}\right)+1\geq 0\,.$
We now express $f(a,b,c)$ in terms of $a$ and $b$. Since c=1abc=1-a-b, we have:
(5) $f(a,b,c)=\dfrac{a}{8}+\dfrac{b}{4}+\dfrac{1-a-b}{2}=\dfrac{1}{2}-\dfrac{3a}{8}-\dfrac{b}{4}\,.$
To relate inequality (4) to $f(a,b,c)$, observe that inequality (4) implies:
$\dfrac{a}{2}+\dfrac{b}{3}\geq-\dfrac{1}{\sqrt{3}}.$
Multiplying both sides by $\dfrac{3}{4}$ given:
$\left(\dfrac{3a}{8}+\dfrac{b}{4}\right)\leq\dfrac{3}{4\sqrt{3}}.$
From Equation (5), we have:
$f(a,b,c)=\dfrac{1}{2}-\left(\dfrac{3a}{8}+\dfrac{b}{4}\right)\,.$
Using the inequality derived above, we obtain:
$f(a,b,c)\leq\dfrac{1}{2}+\dfrac{3}{4\sqrt{3}}=\dfrac{1}{2}+\dfrac{\sqrt{3}}{4}\,.$
By scaling on both x- and y-axes, Lemma 1 implies that using PCHIP interpolation (or any other monotone cubic interpolation for that matter) for splitting buckets in the middle of the gap between two thresholds gives a pair of two buckets with counters smaller than the original bucket counter by at least a constant factor. This technical property is further used in our analysis in Appendix B.
Appendix B Proof of Uniform Error Guarantees and Full Mergeability
We provide a formal analysis of SplineSketch, showing that it guarantees the uniform error (Theorem 3). We first prove that the algorithm can always find a pair of buckets to join when we need to split a bucket (Lemma 2), showing that the algorithm is well-defined. We show both in the most general mergeability setting, i.e., when the sketch is created by an arbitrary sequence of pairwise merge operations executed on single data items (Section B.1). We only assume that all sketches in the process have the same size $k$. The same properties then hold in the streaming setting, which can be simulated by always merging the sketch with a single-item sketch consisting of the buffer only.
In the subsequent analysis, we assume that $k\geq 6$. The precise constant factors in the analysis depend on the interpolation used. Let $\beta>0$ be such that after splitting a bucket, the counters of both new buckets are smaller than the original counter by at least a factor of $1-\beta$. By Lemma 1, we have that $\beta=\frac{1}{2}-\frac{\sqrt{3}}{4}\approx 0.067$ for any monotonic cubic interpolation used to answer queries in the algorithm. For the linear interpolation we have $\beta=0.5$.
B.1. Mergeability setting
For analyzing sketches created by a sequence of pairwise merge operations, it is convenient to define a binary tree $T_{0}$ with $n$ leaves, each representing one data item and inner nodes with two children corresponding to the merge operations. Additionally, we will use inner nodes representing the operations with buckets, namely splitting a bucket or joining a pair of adjacent buckets. That is, we define $T_{0}$ with root representing the final sketch, leaves corresponding the single data items, and inner nodes of three types: (1) merge-nodes with two children, representing merge operations such that the sketch represented by a merge-node is obtained by the union of buckets from the two source sketches represented by the children, as described in Section 3.2, and (2) split-nodes representing one split of a bucket, and (3) join-nodes representing one join of two adjacent buckets; split-nodes and join-nodes have a single child. Each node represents the sketch after the operation is performed. (A technicality is that the number of buckets may possibly be k1k-1 or $k+1$ after a join or split.) Other operations inside the sketch, namely creating initial buckets by equally spaced selection and merging buffer into buckets, need not be represented in $T_{0}$ for our analysis (while equally spaced selection affects the error, it is in total only by $n/k$).
We further label the inner nodes by epochs. Namely, let $n_{0}:=\Theta(k)$ be the initial buffer size and $n_{j}:=1.25\cdot n_{j-1}$ for $j>0$ be the epoch ends. We a call a node $t$ of $T_{0}$ a $j$-node if $j$ is the smallest integer $j\geq 0$ such that the number of items summarized by the sketch represented by $t$ is at most $n_{j}$. Clearly, the labels are non-decreasing on any leaf-to-root path. Moreover, merging two sketches represented by $j$-nodes for $j>0$ results in a sketch represented by a merge-node labeled by $j^{\prime}>j$ as the two source sketches summarize a similar number of items. We thus obtain:
Observation 1.
For any epoch $j>0$, the $j$-nodes of $T_{0}$ form a disjoint union of paths.
The 0-nodes form possibly more complicated subtrees. However, sketches represented by 0-nodes have no buckets and are exact as they store all their items in the buffer.
B.2. Existence of Joinable Pair of Buckets
In describing the consolidate subroutine (Sec. 3.1), we assumed that there is always a pair of buckets that we can join if a bucket counter exceeds the bound (1) or if there is a new minimum or maximum in the buffer. Now we show that it holds.
Lemma 2.
For any node $t$ of the merge tree $T_{0}$, the SplineSketch represented by $t$ contains a joinable pair of adjacent buckets, according to Definition 1.
Proof.
Note that a pair of adjacent buckets is not joinable for one of two reasons: (1) due to bucket bounds, namely that the resulting bucket would have counter exceeding $0.75\cdot C_{b}\cdot n/k$ or (2) due to protected thresholds. Let $n_{t}$ be the number of leaves in the subtree of $t$, i.e., the number of data items summarized by the sketch in $t$. Observe that there are at most $2n_{t}/(0.75\cdot C_{b}\cdot n_{t}/k)=\frac{8}{3}k/C_{b}\leq k/6$ pairs of buckets satisfying case (1), where we use $C_{b}\geq 16$. Thus, the main part of the proof is bounding the number of protected thresholds.
Let $j$ be the label of node $t$. By 1, $j$-nodes in the subtree of $t$ form a path $P_{j}$. If the sketches represented by nodes on $P_{j}$ all have at least $k/3+2$ joinable pairs of buckets, then the lemma holds. Consider the highest node $t^{\prime}$ on the path $P_{j}$ such that the sketch represented by $t^{\prime}$ has at least $k/3+2$ joinable pairs of buckets; we have that $t^{\prime}\neq t$, i.e., $t^{\prime}$ is below $t$, otherwise the lemma holds. Note that the sketch of the lowest node $u$ on $P_{j}$ has at most k1k/6k-1-k/6 pairs that cannot be joined as there are no protected buckets, thus $t^{\prime}$ is well-defined since $k-1-k/6\geq k/3+2$ by $k\geq 6$.
Let $P^{\prime}_{j}$ be the subpath of $P_{j}$ consisting of nodes above $t^{\prime}$. By the definition of $t^{\prime}$, all of them have less than $k/3+2$ joinable pairs of buckets. Thus, the algorithm does not perform splits on sketches on $P^{\prime}_{j}$ due to the heuristic error, that is, it only splits due to exceeding the bucket bound.
To bound the number of split-nodes on $P^{\prime}_{j}$, we consider the following potential of each sketch represented by a node $v$:
(6) $\Psi(v):=\sum_{i=1}^{k}\max\left\{0,b_{i}-0.75\cdot C_{b}\cdot\frac{n_{v}}{k}\right\}\,$,
where $b_{i}$ is the counter of bucket $(\tau_{i-1},\tau_{i}]$ of the sketch of $v$. Clearly, $\Psi(v)\in[0,n_{v}]$ and moreover, the total increase of the potential on path $P^{\prime}_{j}$ is at most $n_{t}$. Next, observe that the potential does not increase by joining buckets (by Def. 1) and that every split at node $v$ of a bucket $i$ with counter $b_{i}>C_{b}\cdot n_{v}/k$ results in two buckets with counters at most $(1-\beta)\cdot b_{i}$. Since $\beta<0.25$ [^1] In case $\beta\geq 0.25$, the constant factors in this proof need to be adjusted., a split thus decreases the potential by at least
$\beta\cdot b_{i}>\beta\cdot C_{b}\cdot\frac{n_{v}}{k}\geq\beta\cdot C_{b}\cdot\frac{n_{t^{\prime}}}{k}\geq\beta\cdot C_{b}\cdot\frac{n_{t}}{1.25\cdot k}$
as every node $v$ on path $P^{\prime}_{j}$ has $t^{\prime}$ in its subtree and $n_{t}\leq 1.25\cdot n_{t^{\prime}}$ by the definition of node labels. Hence, the number of split-nodes on path $P^{\prime}_{j}$ is at most
$\frac{n_{t}}{\beta\cdot C_{b}\cdot\frac{n_{t}}{1.25\cdot k}}=\frac{1.25\cdot k}{\beta\cdot C_{b}}\,.$
We choose $C_{b}$ based on $\beta$ so that the number of split-nodes on path $P^{\prime}_{j}$ is at most $k/18$. As every split results in three protected thresholds, the number of thresholds that get protected on path $P^{\prime}_{j}$ is at most $k/6$. Since there are at least $k/3+2$ joinable pairs of buckets in node $t^{\prime}$, at most $k/6$ of these pair get non-joinable due to falling in case (1), and at most $k/6$ of them get non-joinable due to protected thresholds, there will be a joinable pair of buckets in node $t$. ∎
B.3. Proof of the Uniform Error Guarantee
We now formally prove the worst-case guarantee for SplineSketch, showing that it is fully mergeable [Agarwal et al., 2013, Cormode et al., 2023], i.e., satisfies the error guarantee even if built by an arbitrary sequence of merge operations.
Theorem 3.
Suppose that we build a SplineSketch using any sequence of pairwise merge operations executed on $n$ data items. Then SplineSketch has worst-case rank error of $O(\log(\alpha)\cdot n/k)$.
Proof.
First, we assume w.l.o.g. that there are no items of frequency above $n/k$ by running a Misra-Gries sketch [Misra and Gries, 1982] of size $O(k)$ in parallel, as described in Sec. 3.3. We note that the Misra-Gries sketch is fully mergeable [Agarwal et al., 2013] and that it has error at most $n/k$ for estimating the frequency of any item stored in the sketch. Second, we show an upper bound on the error only for the thresholds of the final sketch. The bound then follows for all other items by the fact that each bucket has $O(n/k)$ items. Thus, in the following we consider a threshold $\tau_{i}$ in the final sketch created by merge operations from $n$ items in an arbitrary way, and analyze the rank error at $\tau_{i}$. We consider the binary tree $T_{0}$ corresponding to merging sketches and operations with buckets described in Section B.1.
The error at threshold $\tau_{i}$ is increased either due to splitting a bucket or due to taking a union of buckets when merging sketches. We first analyze the error from the splits. For the threshold $\tau_{i}$, we consider the following sequence of tree nodes $t_{0},t_{1},t_{2},\dots$ and thresholds $\sigma_{0},\sigma_{1},\sigma_{2},\dots$: Let $\sigma_{0}:=\tau_{i}$ and let $t_{0}$ be the split-node such that the split creates $\tau_{i}$ and there is no node on the path from $t_{0}$ to the root that removes $\tau_{i}$ by a join of buckets. Note that the rank error at $\tau_{i}$ will not increase after the split represented by $t_{0}$, by the description of the algorithm. Then for $a=1,2,\dots$, we define $\sigma_{a}$ and $t_{a}$ as follows: If $\sigma_{a-1}$ is created by a split of a bucket, consider the split of a bucket $(\tau^{\prime},\tau^{\prime\prime}]$ that creates $\sigma_{a-1}=(\tau^{\prime}+\tau^{\prime\prime})/2$ during the consolidate method at $t_{a-1}$. We set $\sigma_{a}:=\tau^{\prime}$ and let $t_{a}$ be the split-node in the subtree of $t_{a-1}$ such that the split at $t_{a}$ creates $\sigma_{a}$ and $t_{a}$ is the closest node to $t_{a-1}$ with the property that on the path from $t_{a}$ to $t_{a-1}$, there is no join of buckets that removes $\sigma_{a}$. Otherwise, $\sigma_{a-1}$ is created by evenly spaced selection of initial buckets at $t_{a-1}$ (i.e., the sketches represented in the subtree of $t_{a-1}$ have no buckets, except for $t_{a-1}$), and we stop this process. Let $m$ be the index when this process stops; that is, at $t_{m}$, threshold $\sigma_{m}$ was created by the evenly spaced selection (Algorithm 1).
To analyze the rank error, for a threshold $\tau$ in the sketch represented by a tree node $t$, we define $\operatorname{Err}^{t}(\tau)$ as the rank error of $\tau$ in the sketch represented by $t$, i.e., the absolute value of the difference between the rank of $\tau$ among items at leaves in the subtree of $t$ (i.e., the items summarized by the sketch at $t$) and the estimated rank of $\tau$, which is the prefix sum of counters of buckets up to $\tau$. We have that $\operatorname{Err}^{t_{m}}(\sigma_{m})=0$, as the evenly space selection is exact at the bucket thresholds, and that $\operatorname{Err}^{t_{0}}(\sigma_{0})=\operatorname{Err}^{t_{0}}(\tau_{i})=\operatorname{Err}^{t_{0}}(r)$, where $r$ is the root, representing the final sketch.
Let $j_{a}$ be the epoch of the sketch represented by $t_{a}$, i.e., $t_{a}$ is a $j_{a}$-node. We claim that for any epoch $j$, there are at most $O(\log(\alpha))$ nodes $t_{a}$ with $j_{a}=j$. Indeed, consider the $j$-nodes on the path $P_{j}$ from the lowest $j$-node to the root of the tree. Since any threshold created by a split in a $j$-node is not removed by a join in a $j$-node due to the protection, any threshold $\sigma_{a}$ created by a $j$-node $t_{a}$ can only be removed by a join of buckets in a $j^{\prime}$-node for $j^{\prime}>j$. Let $b$ be the minimum $a$ such that $j_{a}=j$, i.e., $\sigma_{b}$ is the last threshold created by a $j$-node. Note that after the split in $j$-node $t_{a}$ creating $\sigma_{a}$, the bucket $(\sigma_{a},\sigma^{\prime}_{a}]$ contains $\sigma_{b}$ and as the next threshold $\sigma^{\prime}_{a}$ gets protected, for any $d\in(b,a]$, it holds that $\sigma_{b}\in(\sigma_{d},\sigma^{\prime}_{d}]$ Observe that by the definition of $\alpha$ and since we always split a bucket in the middle, after $O(\log(\alpha))$ splits in $j$-nodes on path $P_{j}$ that involve buckets containing $\sigma_{b}$, the length of the bucket containing $\sigma_{b}$ will be below the smallest distance between distinct items, i.e., the bucket will cover at most one distinct item $x$ of the input. The algorithm may still split the bucket containing $\sigma_{b}$ after the $O(\log(\alpha))$ splits but only $O(1)$ times, because every split decreases the bucket counter by a factor of less than $1-\beta\approx 0.934$, we do not split a bucket with counter below $(C_{b}/100)\cdot n/k$ (even due to the heuristic error), and there are no items of frequency above $n/k$ by the assumption. This proves the claim that there are at most $O(\log(\alpha))$ nodes $t_{a}$ with $j_{a}=j$.
Moreover, each node $t_{a}$ with $j_{a}=j$ increases the error by at most $O(n_{j}/k)$, due to bucket size bounds; specifically for any $a$ with $j_{a}=j$, we have that $\operatorname{Err}^{t_{a}}(\sigma_{a})\leq\operatorname{Err}^{t_{a+1}}(\sigma_{a+1})+O(n_{j}/k)$. Therefore, the total error from splits in $j$-nodes is at most $O(\log\alpha\cdot n_{j}/k)$. Since the values of $n_{j}$ increase geometrically, the total error at $\tau_{i}$ resulting from splits sums to $O(\log\alpha\cdot n/k)$.
To analyze the error from taking the union of buckets in merge-nodes of $T_{0}$, consider a merge-node $t$ such that $\sigma_{a}$ is a threshold of the resulting bucket. Observe that the additional error incurred due to taking the union of buckets is $1/k$ times the number of items summarized by source sketch $S^{\prime}_{t}$ of the merge operation that does not contain $\sigma_{a}$ (if both source sketches contain $\sigma_{a}$, we choose arbitrarily which one is discarded). Further, observe that for any epoch $j$, the total size of such sketches affecting the error at $\sigma_{a}$ with $j_{a}=j$ is at most $O(n_{j})$, since otherwise, the epoch would switch to $j^{\prime}>j$. Since the $n_{j}$’s form a geometric sequence, the total error at $\tau_{i}$ from all merge-nodes is at most $O(n/k)$. Hence, $\operatorname{Err}^{t_{0}}(r)\leq O(\log\alpha\cdot n/k)$. ∎